### v3

# MO-YOLO: End-to-End Multiple-Object Tracking Method with YOLO and Decoder

Abstract
The recent emergence of Transformer-based end-to-end models, exemplified by MOTR, has showcased remarkable multi-object tracking (MOT) performance on datasets like DanceTracker. However, the computational complexities associated with these models pose challenges in both training and deployment. Drawing insights from successful models such as GPT, our proposed MO-YOLO stands out as an efficient and computationally frugal end-to-end MOT solution. MO-YOLO integrates principles from YOLO and RT-DETR, adopting a decoder-centric architecture alongside other complementary structures. By leveraging the RT-DETR decoder and architectural components from YOLOv8, MO-YOLO achieves high-speed performance, shorter training times, and proficient MOT capabilities. On the Dancetrack dataset, MO-YOLO not only achieves parity with MOTR in tracking performance but also surpasses it in terms of speed (MOTR 9.5 FPS, MO-YOLO 19.6 FPS). Moreover, MO-YOLO demonstrates significantly reduced training times and lower hardware requirements compared to MOTR. This research introduces a promising paradigm for efficient end-to-end MOT, highlighting enhanced performance and resource efficiency in a succinct and impactful manner.
Keywords:
Multiple-Object Tracking, Transformer ,YOLO, End-to-End
1 Introduction
Multi-Object Tracking (MOT) is a crucial task in computer vision, aiming to predict object trajectories across consecutive images [2]. The tracking-by-detection approach, relying on object detectors for individual target identification [1, 22, 9], has gained prominence. This method involves associating detections using algorithms like the Kalman filter [3] to form object trajectories. While demonstrating remarkable performance, the intrinsic two-step process—commencing with initial object detection and followed by object tracking—introduces heightened computational complexity and the potential for error propagation.
In recent times, the field has witnessed the emergence of end-to-end MOT models based on Transformer architectures [38, 40, 42], boasting superior performance on datasets such as DanceTracker [29] compared to tracking-by-detection methods. However, the computational demands associated with Transformer models present challenges in terms of both training and deployment, necessitating formidable hardware resources.
In order to address these challenges more effectively, a retrospective analysis of the origins of the Transformer model becomes imperative. The Transformer architecture was initially proposed for the domain of Natural Language Processing (NLP) [31]. As researchers continuously refined this model, the GPT[4] emerged as a prominent highlight in the NLP field over the past few years. GPT, exclusively utilizing the decoder component of the Transformer, achieves pre-training and fine-tuning for various language tasks by stacking multiple decoder layers to construct a deep neural network.
Inspired by the success of GPT, this paper adopts a network structure combining the principles of You Only Look Once (YOLO)[25] and RT-DETR[21], incorporating tracking concepts from MOTR[42]. The resultant model, named MO-YOLO, follows a decoder-based end-to-end MOT approach.
MO-YOLO leverages the decoder from RT-DETR and the block and neck components from YOLOv8 [14], offering advantages such as high speed and shorter training times. This architecture demonstrates proficient performance in multitarget tracking tasks.
This paper pioneers a fusion of concepts from YOLO and the end-to-end MOTR model, resulting in a comprehensive end-to-end MOT model. The contributions of this work can be encapsulated in four fundamental advancements:
  • 1)
    The primary contribution involves the development of a novel end-to-end tracking network. This network amalgamates structural components derived from YOLO, RT-DETR, and the conceptual framework of MOTR, establishing a cohesive and efficient model for Multiple Object Tracking.
  • 2)
    A second key innovation introduces a unique training strategy characterized by a three-stage process. This strategy accelerates model convergence, reducing overall training time and enhancing training process efficiency.
  • 3)
    The third contribution introduces the Tracking Box Selection Process (TBSP) strategy, strategically filtering bounding boxes during training to expedite model convergence and enhance training process efficiency.
  • 4)
    Lastly, we substantiate the feasibility of a decoder based model akin to GPT in the domain of MOT in computer vision.
2 Related work
Figure 1: MOTR’s Streamline the process of updating detect queries and track queries under typical MOT scenarios. The track query set, initially empty, is dynamically adjusted in length. Detect queries play a role in identifying new objects.
2.1 MOTR Series
MOTR [42] is a groundbreaking model that achieves end-to-end multi-object tracking. It extends the concept of object queries from DETR [7] to track queries, allowing for dynamic tracking in video sequences. By employing a learnable positional embedding that interacts with the decoder and feature maps through multi-scale deformable attention, MOTR predicts target positions and classes in each frame. Notably, it establishes implicit temporal associations without the need for explicit data association or post-processing, seamlessly integrating detection and tracking to improve accuracy and robustness. The introduction of Track Query Aggregation with Learned Attention (TALA) is a key innovation, which models long-term temporal relations using a query memory bank and multi-head attention. Additionally, MOTR introduces innovative concepts such as Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to address target conflicts and enhance the integration of temporal information. These design elements collectively position MOTR as a significant milestone in multi-object tracking, elevating both its performance and applicability.
MOTRv2 [47] enhances MOTR’s detection performance by incorporating the pre-trained object detector YOLOX [11] for detection priors. This involves anchoring object queries and utilizing YOLOX-generated proposal boxes as anchors to initialize the object queries. This strategy not only improves detection but also mitigates conflicts between detection and association tasks while preserving query propagation characteristics.
MeMOTR [10] introduces track queries, modeling target trajectories throughout the video and predicting them iteratively in each frame. It implements a tracklet-aware label assignment method to achieve one-to-one assignment between track queries and target trajectories, considering newborn object queries.
MOTRv3 [40] and CO-MOT [37] share a common focus on addressing unfair label assignment between detection and tracking queries in MOTR, aiming to balance training samples and improve end-to-end multi-object tracking performance. MOTRv3 employs a release-fetch supervision strategy for balanced label assignment, while CO-MOT uses a coopetition label assignment strategy to increase positive samples for detection queries and introduces "shadows" to enrich query quantity and diversity.
2.2 YOLO
YOLO[25] has transformed object detection into a regression model, achieving efficient target detection. It partitions the image into grids, with each grid responsible for predicting the target box and class information within its corresponding region. The object detection task is treated as a regression problem for coordinates and dimensions. This unique approach simplifies the detection process, enabling YOLO to simultaneously process multiple targets in a single forward pass, thus achieving outstanding speed and efficiency. The concept of this regression model provides a solid foundation for real-time object detection.
Due to its relatively high efficiency and accuracy, the SORT[3] initially applied YOLO in the field of MOT. However, with the proliferation of such models, YOLO has become widely used in the MOT domain. Nevertheless, this approach still introduces additional processing layers, potentially leading to some errors. Therefore, in this paper, we construct an end-to-end tracking model based on the YOLO framework.
2.3 RT-DETR
DETR [50] revolutionizes object detection by introducing the Transformer architecture, transforming the task into a multi-class classification problem. It excels in precise object localization and modeling object relationships, making it particularly suitable for applications like multi-object tracking. Despite its effectiveness with limited annotated data, DETR comes with the drawback of high computational complexity, demanding significant computational resources.
RT-DETR [21] addresses DETR’s computational challenges, presenting the first real-time usable DETR model. It tackles complexity with an efficient hybrid encoder that manages multi-scale features through smart intra-scale interactions and cross-scale fusion decoupling. This significantly reduces computational costs, achieving real-time object detection.
RT-DETR’s decoder module incorporates deformable attention, aligning with attention mechanisms in various DETR models [50, 36, 51]. By dynamically generating attention sampling points, it adapts to diverse feature distributions and target shapes, enhancing adaptability. Unlike some DETR models [36, 43, 8], RT-DETR skips the use of a reconstruction network, relying on the decoder’s output for final predictions. This simplifies the architecture while maintaining efficiency and performance.
3 MO-YOLO
Figure 2: The overall structure of MO-YOLO consists of an architecture that combines YOLOv8’s backbone and neck, facilitating the extraction of image features for each frame. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. When the model is trained using the TALA strategy, the TBSP can be skipped.
3.1 MO-YOLO Architecture
MO-YOLO’s overall architecture is illustrated in the provided Fig.2. In this setup, a video sequence is input to YOLOv8’s backbone, neck to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the RT-DETR decoder. The learning pattern of $q_{d}$ is similar to RT-DETR[21].
For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder. These queries interact with the image features in the decoder to generate hidden states for bounding box predictions. Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
3.2 Tracking Box Selection Process
When training a detection model using the detection method and subsequently employing an end-to-end tracking approach, we observed a gradual increase in the number of tracking boxes for the same object over time. Upon examination, this phenomenon arises from the fact that each newly generated detect query is consistently assigned to an already tracked object. This is due to the nature of DETR, where a detect query can potentially be assigned to any object in the image, as label assignment is determined through binary matching of all detect queries and ground truth.
In MOTR[42], a solution to this issue is introduced through the utilization of TALA. This strategy gradually suppresses the generation of detect queries similar to track queries during the training process, effectively addressing the identified problem. However, this process increases the number of parameters that the network needs to automatically adjust, resulting in a prolonged convergence time for the model. Therefore, this paper proposes a simple yet effective approach named tracking box selection process (TBSP). Experimental results demonstrate the efficacy of TBSP, revealing that this method allows the model to achieve comparable performance to fully trained models even without extensive training. Additionally, the simplicity of TBSP ensures that it introduces minimal additional computational burden.
Figure 3: Refinement process for MO-YOLO on partially trained track queries. When MO-YOLO undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of MO-YOLO aligns with MOTR, as illustrated in the Fig. 1.
We know that both track query and detect query contain information about the object’s bounding box(BBox). Let’s denote the BBox in $\hat{\omega}_{tr}^{t-1}$ as $[BBox_{1}^{tr}$, $BBox_{2}^{tr},...,BBox_{n}^{tr}]$, where $n$ is the number of track queries at time $t-1$. Each query in $\hat{\omega}_{tr}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
For each element in $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$, calculate the Intersection over Union (IOU) with each element in $[BBox_{1}^{tr},BBox_{2}^{tr},...,BBox_{n}^{tr}]$. If the IOU is greater than $\delta_{iou}$, the BBox corresponding to that detect Query will be discarded. Finally, the remaining detect queries are denoted as $\omega_{\det_{f}}^{t-1}$.
Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2.
Similarly, assuming the track query at frame $t$ (where $t>2$) is denoted as $\omega_{tr}^{t}$, it can be expressed as:
$\hat{\omega}_{tr}^{t}\,\,=\omega_{det_{f}}^{t-1}\cup\,\,\hat{\omega}_{tr}^{t-1}\,\,$ (1)
The feasibility of TBSP is illustrated in the Fig. 4. Incidentally, although TBSP makes MO-YOLO appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our model compared to MOTR in CO-MOT[38] and MOTRv3[40], we have not adopted the labeling strategy used in CO-MOT and MOTRv3.
3.3 Decoder Embeddings and Query Positions
Differentiating itself from MOTR, MO-YOLO’s decoder employs a methodology akin to RT-DETR for its embeddings. In contrast to MOTR, MO-YOLO’s decoder input doesn’t originate from a distinct encoder; rather, it exploits RT-DETR’s IoU-aware query selection mechanism. Building upon prior works such as [35, 43, 50, 16], this process strategically selects high-quality image features as the initial object embeddings for the decoder, emphasizing IoU scores.
It’s essential to clarify that in this context, the term "embeddings" specifically pertains to the initial embeddings of detection queries. The subsequent embeddings of track queries undergo generation through a TAN network, aligning with MOTR’s established process.
To expedite model convergence, we implement a Query Positions generation strategy akin to CO-MOT. Drawing inspiration from DAB-DETR [19], we employ a fixed transformation to convert reference points into Query Positions. While experimentation with the RT-DETR approach involving a Multilayer Perceptron (MLP) to transform reference points shows no significant performance discrepancy, it is important to note that the use of MLP substantially increases training time. By the way, it is worth mentioning that this process differs significantly from MOTR. In MOTR, Query Positions are generated using a set of learnable parameters, whereas in MO-YOLO, they are generated using reference points.
3.4 Loss and Training Methods
Due to the unique training mode adopted by MOTR, its batch size in a single GPU is constrained to 1. While this training approach enables the model to effectively learn the temporal information of the targets, it concurrently requires training the network to extract features of target appearances. To expedite the model training speed, a three-stage training strategy is employed in this study. In the first stage, the emphasis is placed on training the network to extract features of target appearances; the second stage focuses on learning temporal information; and the final stage comprehensively enhances the model’s performance by integrating the learning of temporal information and appearance features. This strategy is implemented to overcome the constraints posed by the unique training mode of MOTR, aiming to achieve more efficient model training.
In the initial stage, the network is trained as a detection network, and the loss function is formulated defined as Eq.2:
$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)}{V_{i}}$ (2)
$\mathcal{L}$ is the loss of single frame, which is similar to the detection loss in DETR, and it is shown in Eq.3.
$\mathcal{L}\left(\left.\widehat{Y}_{t}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}$ (3)
where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[17], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[26], $\lambda_{cls},\lambda_{l1},\lambda_{giou}$ are corresponding weight coefficients.
In this stage, the network is capable of initially learning the appearance features of objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Additionally, various data augmentation techniques [44, 13, 32] can be employed to expedite the model’s extraction of appearance features.
The second stage of training is conducted using MOTR’s Collective Average Loss. During this phase, only MOTR’s TAN and our TBSP are utilized. The loss functions used in this stage and the third stage are as Eq.4. Simultaneously, it is important to note that, at this juncture, the determination of the newly generated track queries relies solely on the scores of object queries and whether they have been filtered out by TBSP, as more advanced strategies like TALA have not been employed.
$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\sum_{n=1}^{N}{\left(\mathcal{L}\left(\left.\widehat{Y}_{tr}^{t}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)+\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)\right)}}{\sum_{n=1}^{N}{\left(V_{t}\right)}},N=5$ (4)
where $V_{i}=V_{tr}^{i}+V_{det}^{i}$ denotes the total number of ground-truths objects at frame $i$. $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively.
The third stage of training involves the utilization of TALA policy during the training phase, while simultaneously deactivating the TBSP. Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance. This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, the TBSP policy can be turned off during inference with negligible impact on overall performance.
In addition, it is important to note that while the first and second stages of MO-YOLO share similarities in generating a new track query from the object query, discrepancies may arise in the third stage. Tracking objects with the track query does not involve additional techniques, and the process across the three stages of MO-YOLO closely resembles that of MOTR.
4 Experiments
4.1 Datasets and Metrics
This paper introduces MO-YOLO, an innovative MOT method proficient in addressing intricate scenarios involving occlusions, interactions, and diverse motions. To validate its performance, we evaluate the algorithm on three demanding datasets: DanceTrack[29], MOT17[24], and KITTI[12].
DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures. Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
MOT17 focuses on multiple object tracking in public spaces, primarily pedestrians. The dataset includes seven scenes, divided into training and testing clips, providing object detection annotations suitable for both online and offline tracking approaches.
KITTI serves as a research dataset for autonomous driving, encompassing stereo, optical flow, odometry, 3D object detection, and tracking tasks. Captured across diverse environments using advanced sensors, it provides raw data and task-specific benchmarks.
Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics. The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[27, 20]. These metrics collectively provided a comprehensive and robust evaluation of our model’s tracking performance. HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks. AssA measured the quality of associations between objects, while IDF1 quantified the accuracy of identity tracking. MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches. These metrics formed the basis for a thorough evaluation of our model’s tracking capabilities.
4.2 Implementation Details
The rationale for adopting MO-YOLO originates from our observation of the high training costs and suboptimal inference speed associated with MOTR. Thus, our aim is to propose a new baseline model surpassing MOTR’s performance. To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [10], MOTRv2 [47], MOTRv3 [40], and CO-MOT [38]. To achieve a comparable parameter count with MOTR, we simultaneously utilize YOLOv8-L [14] and RT-DETR-L [21] in constructing MO-YOLO (henceforth referred to as MO-YOLO-L concerning model size). The initial weights are obtained from a model pre-trained on the COCO dataset [18].
The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU. For equitable comparison with MOTR, training on the MOT17 dataset was performed on a single NVIDIA GeForce RTX 2080ti. Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU. One more thing, he image size for the input network is $640\times 640$.
During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1. For MOT17, like MOTR, we integrated extra data from the CrowdedHuman dataset [28], employing random displacement to generate video segments with pseudo-trajectories. Training comprised 120 epochs in the first stage, followed by 30 epochs in the second. The third stage extended to 20 epochs with TBSP, or 55 without. For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 6, and the third, if TBSP was used, was 6 epochs, or 15 without TBSP. For KITTI, initial training involved 80 epochs, followed by 25 in the second stage. The third stage consisted of 30 epochs with TBSP or 70 without. Additionally, approximately 5k images from BDD100k [41] were selected for additional training.
Note: With the exception of Table 4(b), all data concerning MO-YOLO in this paper pertains to MO-YOLO-L. Additionally, unless stated otherwise, the provided data pertains to the final end-to-end model.
4.3 Performance Evaluation and Comparison
In Table 1, a comprehensive comparison of training time and speed was conducted between MO-YOLO and the MOTR series.
(a) Comparison of Training Time on MOT17
Method Device Training Time
MOTR[42] 8 2080ti $\sim$ 96.0 hours
MO-YOLO stage1 1 2080ti 12.1 hours
MO-YOLO stage2 1 2080ti 30.8 hours
MO-YOLO(TBSP) 1 2080ti 43.4 hours
MO-YOLO 1 2080ti 68.7 hours
(b) Comparison of Speed on Dancetrack
Method FPS
MOTR[42] 9.5
MOTRv2[47] 6.9
MOTRv3[40] 10.6
MO-YOLO(TBSP) 18.5
MO-YOLO 19.6
Table 1: Comparison of training time (including all previous stages) and speed between MO-YOLO and MOTR series on a single V100 GPU. The data for the MOTR series is sourced from [42], [40] or their open source address.
It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations. This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
MO-YOLO demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed. The primary reason behind this efficiency gain lies in the architectural simplicity of MO-YOLO. Unlike MOTR, MO-YOLO adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design. This reduction in complexity enables MO-YOLO to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU. In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training. The absence of an encoder in MO-YOLO contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
On the Dancetrack test dataset, MO-YOLO(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3). Particularly noteworthy is the remarkable 19.6 FPS achieved by MO-YOLO, emphasizing its outstanding performance in inference speed.
Table 2: Performance comparison between MO-YOLO and existing methods on the Dancetrack[29] dataset under the private detection protocols. Numbers are highlighted in bold when they correspond to either MO-YOLO or the superior metric from MOTR.
Methods HOTA AssA DetA IDF1 MOTA
Non-End-to-end:
ByteTrack[45] 47.3 31.3 71.6 52.5 89.5
GTR[49] 48.0 31.9 72.5 50.3 84.7
DST-track[6] 51.9 34.6 82.3 51.0 84.9
OC-SORT[5] 54.6 38.0 80.4 54.2 89.4
End-to-end:
MeMOTR[10] 63.4 52.3 77.0 65.5 85.4
MOTR[42] 54.2 40.2 73.5 51.5 79.7
MO-YOLO(Ours) 52.0 35.7 75.9 51.4 83.5
Table 2, Table 3, and Table 4 exhibit the final results of end-to-end MO-YOLO on the Dancetrack, MOT17, and KITTI datasets, respectively. In contrast to MOTR, MO-YOLO demonstrates competitive performance across multiple metrics on both datasets. It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3. The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our MO-YOLO. MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training. Consequently, comparing its tracking performance directly with that of MO-YOLO and MOTR may not be entirely fair.
Table 3: Performance comparison between MO-YOLO and existing methods on the MOT17[24] dataset.
method HOTA AssA DetA IDF1 MOTA
Non-End-to-end:
FairMOT[46] 59.3 58.0 60.9 72.3 73.7
ByteTrack[45] 62.8 62.2 63.8 77.2 78.9
StrongSORT[49] 63.5 63.7 63.6 78.3 78.5
UTM[39] 64.0 62.5 65.9 78.7 81.8
OC-SORT[5] 63.2 63.2 63.2 77.5 78.0
Deep OC-SORT[22] 64.9 65.9 64.1 80.6 79.4
End-to-end:
MeMOTR[10] 58.8 58.4 59.6 69.0 72.5
MOTR[42] 57.2 55.8 58.9 59.4 71.9
MO-YOLO(Ours) 55.5 55.9 55.8 70.1 66.8
For Dancetrack (Table 2), MO-YOLO achieves comparable results to MOTR, with slightly lower scores in HOTA and AssA but outperforming MOTR in DetA. Specifically, MO-YOLO achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
In the MOT17 (Table 3), MO-YOLO demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA. Despite trailing slightly behind MOTR in these metrics, MO-YOLO compensates with a comparable DetA and excels in IDF1 by almost 10 points. This advantage can be attributed to MO-YOLO’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand. Additionally, MO-YOLO maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
Table 4: Performance comparison between MO-YOLO and existing methods on the KITTI[12] dataset.
Car Pedestrian
Tracker HOTA MOTA AssA HOTA MOTA AssA
Non-End-to-end:
CenterTr[48] 73.0 88.8 71.1 40.3 53.8 36.9
PermaTr[30] 77.4 90.8 77.6 47.4 65.0 43.6
PolarMOT[15] 75.1 80.0 76.9 43.6 46.9 48.1
TripletTrack[23] 73.6 84.3 74.7 50.0 46.5 57.8
End-to-end:
MO-YOLO 72.2 83.2 73.8 51.5 56.8 58.4
Table 4 provides an overview of MO-YOLO’s performance on the KITTI dataset. MO-YOLO achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively. These results position MO-YOLO as a strong contender in multi-object tracking on the KITTI.
In conclusion, MO-YOLO stands out as a highly advantageous model in the realm of multi-object tracking, demonstrating superior object detection accuracy and overall tracking performance compared to MOTR, as evidenced by competitive results on diverse datasets, including Dancetrack, MOT17, and KITTI. The notable improvements observed in MeMOTR relative to MOTR suggest that MO-YOLO has significant untapped potential for performance enhancement. Positioned as a robust alternative, MO-YOLO can serve as an excellent new baseline, driving advancements in end-to-end models within the field of multi-object tracking.
4.4 Ablation Study
To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack dataset. The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set. Additionally, the impact of two methods for generating Query Positions—utilizing the formula from DAB-DETR[19] or employing an MLP—on the model’s performance was investigated. Notably, MO-YOLO-L, MO-YOLO-M, and MO-YOLO-S differ based on their foundations in YOLOv8-L,YOLOv8-M, and YOLOv8-S, respectively, providing insights into model variations and their implications.
(a) Performance of MO-YOLO at different training stages Training Stage HOTA AssA MOTA Initial 38.5 21.6 76.8 Intermediate 44.5 25.5 84.2 Final (TBSP) 51.9 34.6 84.9 Final 53.6 35.9 84.7 (b) Performance across different model sizes Model HOTA MOTA FPS MO-YOLO-L 53.6 84.7 19.6 MO-YOLO-M 49.7 82.9 22.7 MO-YOLO-S 42.9 79.3 26.2
(c) The impact of $\delta_{iou}$ on model performance when using TBSP $\delta_{iou}$ HOTA AssA MOTA 0.3 49.7 32.3 79.1 0.5 51.9 34.6 84.9 0.7 51.3 34.8 83.2 0.9 50.0 32.8 81.3 (d) Effect of query position generation strategy Strategy HOTA MOTA FPS MLP 53.3 85.1 18.5 Formula 53.6 84.7 19.6
Table 5: Ablation Studies on Our Proposed MO-YOLO on the DanceTrack validation set.
The performance of MO-YOLO improves consistently across different training stages, peaking in the final stage. Different model sizes (MO-YOLO-L, MO-YOLO-M, MO-YOLO-S) achieve a balance between accuracy and computational efficiency. The similarity in their speeds is likely attributable to the consistent scale of the decoder. The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance. Position generation strategies, MLP and Formula, yield similar performance in terms of detection accuracy, with slight differences in FPS. Overall, the ablation studies demonstrate the effectiveness of MO-YOLO across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
Figure 5: MO-YOLO and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 dataset.
5 Conclusion and Limitation
This paper presents MO-YOLO, an efficient and lightweight end-to-end multi-object tracking model that optimizes computational resources. By integrating YOLO and RT-DETR principles, the study establishes a novel tracking paradigm with accelerated training strategies. While MO-YOLO outperforms MOTR in resource efficiency, it marginally lags behind in performance, attributed to factors like YOLO’s detection capabilities and reduced input image dimensions ($640\times 640$). Despite enhancing computational efficiency, the dimension reduction raises concerns about potential information loss, impacting overall effectiveness. Additionally, latent issues from MOTR may surface in MO-YOLO, as depicted in Fig.5. This research contributes to advancing real-time computer vision applications.
Appendix 0.A Appendix
0.A.1 Disclaimer
The data used in this paper is sourced from publicly available datasets related to human subjects. While every effort has been made to ensure the accuracy and reliability of the data, the authors make no representations or warranties of any kind, express or implied, about the completeness, accuracy, reliability, suitability, or availability with respect to the data. The authors shall not be liable for any loss or damage, including but not limited to indirect or consequential loss or damage, arising from the use of the data in this research paper. Users of the data are encouraged to independently verify the accuracy and reliability of the data before making any decisions or taking any actions based on it.
0.A.2 More Comprehensive Analysis of Experimental Results and Model Parameter
The initial motivation for building MO-YOLO was not because of its slow inference speed, but rather because of its high training resource requirements. With 8 V100 GPUs using pre-trained weights, it took approximately 84 hours to train 200 epochs on an extended MOT17[24] with only about 10,000 images. Each epoch took around 0.42 hours, and on average, each GPU could train on approximately 3,000 images per hour.
It is worth noting that MOTR[42] is based on Deformable-DETR[50], which was trained on the MSCOCO dataset[18] with approximately 110,000 images. In single-scale mode (which can be considered similar to MOTR’s training), using eight V100 GPUs, each epoch took approximately 0.4 hours, and on average, each GPU could train on around 30,000 images per hour.
Upon careful consideration of the reasons, the first aspect to be contemplated is the batch size. The collective average loss (CAL) mechanism and the tracklet-aware label assignment (TALA) strategy employed in MOTR enable the network to learn some latent temporal information, consequently restricting the batch size on a single card to 1. Setting the batch size to 1 may result in underutilization of the corresponding computational resources (e.g., GPU, TPU, or other possible devices). Additionally, MOTR is trained in a one-shot manner, requiring the model not only to grasp motion information but also to learn object appearance features. Hence, we proposed a multi-step training strategy to first allow the model to learn appearance features before proceeding to learn object motion characteristics. Experimental results confirm the effectiveness of this strategy.
As for the TBSP, its initial purpose was simply to provide the detection network with a basic tracking capability. However, it also pleasantly surprised us by further reducing training time. Here, we also provide the required rounds and results if TALA were to be used directly without the TBSP strategy, as shown in Table 6. The training here is conducted on a single NVIDIA 4090 GPU. And this training process is not included in the first stage
Table 6: Comparison of MO-YOLO with and without using TBSP on the Dancetrack val dataset.
Initial Intermediate Final HOTA MOTA Epochs Total Time
53.6 84.7 27 56.7 hours
52.9 83.8 40 85.2 hours
In order to provide a more comprehensive illustration of the characteristics of our model, we list the sizes of different models, namely MO-YOLO-L (45.8M), MO-YOLO-M (29.9M), and MO-YOLO-S (19.7M), as depicted in Fig. 6.
Figure 6: The model sizes of MOTR and MO-YOLO, as well as their FPS and HOTA on the Dancetrack.
The table compares the performance of MO-YOLO with and without using the TBSP strategy on the Dancetrack validation dataset across different training phases: Initial, Intermediate, and Final. Comparison of the two rows reveals that employing TBSP throughout all phases leads to slightly better performance in terms of HOTA and MOTA, with a significantly lower total training time. This underscores the importance of the TBSP strategy in reducing the overall training duration while maintaining or even improving tracking accuracy.
We hypothesize that the efficacy of the TBSP strategy may stem from its endowment of the model with a degree of autonomy, facilitating the preliminary acquisition of object motion information that would otherwise necessitate numerous training iterations. Analogously, the strategy resembles the use of training wheels in bicycle riding, which assist novices in achieving smooth operation while providing an opportunity to experience the sensation of riding. Consequently, this approach aids in the comprehensive mastery of the skill. We posit that a comparable scenario may manifest within this context.
0.A.3 RT-MOTR
At the same time, to demonstrate the advantages of our MO-YOLO models constructed using YOLO and RT-DETR relative to end-to-end tracking models built directly on RT-DETR, we have also constructed such an end-to-end tracking model based on RT-DETR (referred to as RT-MOTR), and conducted performance testing on it.
NOTE: in Table 7, both the train time and epochs refer to the second and third stages like Table 6. For the first stage, both models were trained for 15 epochs, and the training device is a single 4090, while the testing device is a single V100.
The table compares the performance of MO-YOLO and RT-MOTR on the Dancetrack validation dataset. It presents various metrics including HOTA (Higher Order Tracking Accuracy), training epochs, training speed (hours per epoch), total training time, and frames per second (FPS).
MO-YOLO achieves a HOTA of 53.6 with 27 training epochs, an average training speed of 2.08 hours per epoch, a total training time of 56.7 hours, and an FPS of 19.6.
In comparison, RT-MOTR achieves a HOTA of 53.4 with 35 training epochs, an average training speed of 2.27 hours per epoch, a total training time of 90.8 hours, and an FPS of 20.3.
Comparing the performance metrics of the two methods, we observe that MO-YOLO slightly outperforms RT-MOTR in terms of HOTA, while also having a shorter training time albeit with a slightly lower FPS. This indicates that MO-YOLO exhibits advantages in both tracking accuracy and training efficiency.
Table 7: Comparison of MO-YOLO and RT-MOTR on the Dancetrack val dataset.
Methods HOTA Train Epochs Speed (hours/epoch) Train Times FPS
MO-YOLO 53.6 27 2.08 56.7 h 19.6
RT-MOTR 53.4 35 2.27 79.5 h 20.3
Regarding why the speeds of these two models are comparable, our experimental findings suggest that the primary time-consuming component is the decoder responsible for tracking. Therefore, optimizing the decoder represents a crucial area for future investigation.
This outcome underscores the rationale behind integrating elements from YOLOv8 and RT-DETR to construct a rapid end-to-end tracking network. Furthermore, the observed disparity in training efficiency between MO-YOLO and RT-DETR can be attributed to the former’s notably simpler architecture.
Moreover, our decision to adopt YOLO as the backbone of the network is motivated by its propensity for swifter updates compared to DETR-based models. Recent developments such as YOLOv9[34] exemplify this trend. While YOLOv9 may trade off training time for enhanced precision and speed (resulting in at least a twofold increase in training duration compared to YOLOv7[33]), we anticipate future iterations will offer refined improvements. Such advancements will, in turn, facilitate further enhancements to MO-YOLO.

### v4

@@
-MO-YOLO: End-to-End Multiple-Object Tracking Method with YOLO and Decoder
+DecoderTracker: Decoder-Only Method for Multiple-Object Tracking
 Abstract
-The recent emergence of Transformer-based end-to-end models, exemplified by MOTR, has showcased remarkable multi-object tracking (MOT) performance on datasets like DanceTracker.
-However, the computational complexities associated with these models pose challenges in both training and deployment.
-Drawing insights from successful models such as GPT, our proposed MO-YOLO stands out as an efficient and computationally frugal end-to-end MOT solution.
-MO-YOLO integrates principles from YOLO and RT-DETR, adopting a decoder-centric architecture alongside other complementary structures.
-By leveraging the RT-DETR decoder and architectural components from YOLOv8, MO-YOLO achieves high-speed performance, shorter training times, and proficient MOT capabilities.
-On the Dancetrack dataset, MO-YOLO not only achieves parity with MOTR in tracking performance but also surpasses it in terms of speed (MOTR 9.5 FPS, MO-YOLO 19.6 FPS).
-Moreover, MO-YOLO demonstrates significantly reduced training times and lower hardware requirements compared to MOTR.
-This research introduces a promising paradigm for efficient end-to-end MOT, highlighting enhanced performance and resource efficiency in a succinct and impactful manner.
-Keywords:
-Multiple-Object Tracking, Transformer ,YOLO, End-to-End
+Decoder-only models, such as GPT[1], have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer models.
+Over the years, end-to-end models based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking.
+However, the significant computational resource consumption of these models leads to less friendly inference speeds and training times.
+To address these issues, this paper attempts to construct a lightweight Decoder-only model: DecoderTracker for end-to-end multi-object tracking.
+Specifically, drawing on some real-time detection models, we have developed an image feature extraction network which can efficiently extract features from images to replace the encoder structure.
+In addition to minor innovations in the network, we analyze the potential reasons for the slow training of MOTR-like models and propose an effective training strategy to mitigate the issue of prolonged training times.
+On the DanceTrack dataset, without any bells and whistles, DecoderTracker’s tracking performance slightly surpasses that of MOTR, with approximately twice the inference speed.
+Furthermore, DecoderTracker requires significantly less training time compared to MOTR.
 1 Introduction
-Multi-Object Tracking (MOT) is a crucial task in computer vision, aiming to predict object trajectories across consecutive images [2].
-The tracking-by-detection approach, relying on object detectors for individual target identification [1, 22, 9], has gained prominence.
-This method involves associating detections using algorithms like the Kalman filter [3] to form object trajectories.
-While demonstrating remarkable performance, the intrinsic two-step process—commencing with initial object detection and followed by object tracking—introduces heightened computational complexity and the potential for error propagation.
-In recent times, the field has witnessed the emergence of end-to-end MOT models based on Transformer architectures [38, 40, 42], boasting superior performance on datasets such as DanceTracker [29] compared to tracking-by-detection methods.
-However, the computational demands associated with Transformer models present challenges in terms of both training and deployment, necessitating formidable hardware resources.
-In order to address these challenges more effectively, a retrospective analysis of the origins of the Transformer model becomes imperative.
-The Transformer architecture was initially proposed for the domain of Natural Language Processing (NLP) [31].
-As researchers continuously refined this model, the GPT[4] emerged as a prominent highlight in the NLP field over the past few years.
-GPT, exclusively utilizing the decoder component of the Transformer, achieves pre-training and fine-tuning for various language tasks by stacking multiple decoder layers to construct a deep neural network.
-Inspired by the success of GPT, this paper adopts a network structure combining the principles of You Only Look Once (YOLO)[25] and RT-DETR[21], incorporating tracking concepts from MOTR[42].
-The resultant model, named MO-YOLO, follows a decoder-based end-to-end MOT approach.
-MO-YOLO leverages the decoder from RT-DETR and the block and neck components from YOLOv8 [14], offering advantages such as high speed and shorter training times.
-This architecture demonstrates proficient performance in multitarget tracking tasks.
-This paper pioneers a fusion of concepts from YOLO and the end-to-end MOTR model, resulting in a comprehensive end-to-end MOT model.
-The contributions of this work can be encapsulated in four fundamental advancements:
+Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images.
+The success of transformer[2] models in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT models like MOTR[3] have been introduced in the MOT domain.
+These models offer advantages over the widely used tracking-by-detection methods[4, 5, 6, 7], such as eliminating the need for manual design or selection of features.
+Additionally, some enhanced methods[8, 9, 10] of MOTR have demonstrated the potential of this model class in terms of tracking accuracy.
+However, the inference speed of these models is often suboptimal, and their RNN[11] like linear inference structure makes them less GPU friendly during training.
+[…]
+At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the model to engage in less stringent learning might also be effective.
+Is there a possibility that this process could yield unexpected benefits?
+Based on this idea, we proposed the Tracking Box Selection Process (TBSP).
+Although the methods we propose have not yet enabled the inference speed of MOTR-like models to achieve real-time performance, we believe that this paper could provide valuable insights for future researchers.
+In summary, the contributions of this paper are as follows:
   • 1)
-    The primary contribution involves the development of a novel end-to-end tracking network.
-This network amalgamates structural components derived from YOLO, RT-DETR, and the conceptual framework of MOTR, establishing a cohesive and efficient model for Multiple Object Tracking.
+    The primary contribution involves the development of a novel end-to-end tracking network, DecoderTracker, which is a decoder-only method.
+Compared to MOTR, this network achieves faster inference speeds and improved tracking performance.
   • 2)
@@
   • 3)
-    The third contribution introduces the Tracking Box Selection Process (TBSP) strategy, strategically filtering bounding boxes during training to expedite model convergence and enhance training process efficiency.
-  • 4)
-    Lastly, we substantiate the feasibility of a decoder based model akin to GPT in the domain of MOT in computer vision.
+    The third contribution introduces TBSP, a strategy that can serve as a preliminary weakly supervised training approach for end-to-end MOT model training.
 2 Related work
-Figure 1: MOTR’s Streamline the process of updating detect queries and track queries under typical MOT scenarios.
+Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios.
 The track query set, initially empty, is dynamically adjusted in length.
@@
 2.1 MOTR Series
-MOTR [42] is a groundbreaking model that achieves end-to-end multi-object tracking.
-It extends the concept of object queries from DETR [7] to track queries, allowing for dynamic tracking in video sequences.
-By employing a learnable positional embedding that interacts with the decoder and feature maps through multi-scale deformable attention, MOTR predicts target positions and classes in each frame.
-Notably, it establishes implicit temporal associations without the need for explicit data association or post-processing, seamlessly integrating detection and tracking to improve accuracy and robustness.
-The introduction of Track Query Aggregation with Learned Attention (TALA) is a key innovation, which models long-term temporal relations using a query memory bank and multi-head attention.
-Additionally, MOTR introduces innovative concepts such as Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to address target conflicts and enhance the integration of temporal information.
-These design elements collectively position MOTR as a significant milestone in multi-object tracking, elevating both its performance and applicability.
-MOTRv2 [47] enhances MOTR’s detection performance by incorporating the pre-trained object detector YOLOX [11] for detection priors.
-This involves anchoring object queries and utilizing YOLOX-generated proposal boxes as anchors to initialize the object queries.
-This strategy not only improves detection but also mitigates conflicts between detection and association tasks while preserving query propagation characteristics.
-MeMOTR [10] introduces track queries, modeling target trajectories throughout the video and predicting them iteratively in each frame.
-It implements a tracklet-aware label assignment method to achieve one-to-one assignment between track queries and target trajectories, considering newborn object queries.
-MOTRv3 [40] and CO-MOT [37] share a common focus on addressing unfair label assignment between detection and tracking queries in MOTR, aiming to balance training samples and improve end-to-end multi-object tracking performance.
-MOTRv3 employs a release-fetch supervision strategy for balanced label assignment, while CO-MOT uses a coopetition label assignment strategy to increase positive samples for detection queries and introduces "shadows" to enrich query quantity and diversity.
-2.2 YOLO
-YOLO[25] has transformed object detection into a regression model, achieving efficient target detection.
-It partitions the image into grids, with each grid responsible for predicting the target box and class information within its corresponding region.
-The object detection task is treated as a regression problem for coordinates and dimensions.
-This unique approach simplifies the detection process, enabling YOLO to simultaneously process multiple targets in a single forward pass, thus achieving outstanding speed and efficiency.
-The concept of this regression model provides a solid foundation for real-time object detection.
-Due to its relatively high efficiency and accuracy, the SORT[3] initially applied YOLO in the field of MOT.
-However, with the proliferation of such models, YOLO has become widely used in the MOT domain.
-Nevertheless, this approach still introduces additional processing layers, potentially leading to some errors.
-Therefore, in this paper, we construct an end-to-end tracking model based on the YOLO framework.
-2.3 RT-DETR
-DETR [50] revolutionizes object detection by introducing the Transformer architecture, transforming the task into a multi-class classification problem.
-It excels in precise object localization and modeling object relationships, making it particularly suitable for applications like multi-object tracking.
-Despite its effectiveness with limited annotated data, DETR comes with the drawback of high computational complexity, demanding significant computational resources.
-RT-DETR [21] addresses DETR’s computational challenges, presenting the first real-time usable DETR model.
-It tackles complexity with an efficient hybrid encoder that manages multi-scale features through smart intra-scale interactions and cross-scale fusion decoupling.
-This significantly reduces computational costs, achieving real-time object detection.
-RT-DETR’s decoder module incorporates deformable attention, aligning with attention mechanisms in various DETR models [50, 36, 51].
-By dynamically generating attention sampling points, it adapts to diverse feature distributions and target shapes, enhancing adaptability.
-Unlike some DETR models [36, 43, 8], RT-DETR skips the use of a reconstruction network, relying on the decoder’s output for final predictions.
-This simplifies the architecture while maintaining efficiency and performance.
-3 MO-YOLO
-Figure 2: The overall structure of MO-YOLO consists of an architecture that combines YOLOv8’s backbone and neck, facilitating the extraction of image features for each frame.
+Unlike tracking-by-detection models exemplified by SORT[4, 14, 5], and previous transformer-based MOT models[15, 16], MOTR[3] is the first model to achieve end-to-end multi-object tracking.
+It extends the object query from DETR[17] into a track query, thereby enabling tracking capabilities.
+It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing.
+Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for modeling long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration.
+These features establish MOTR as a significant advancement in multi-object tracking technology.
+MOTRv2[18] enhances the detection capabilities of the original MOTR by integrating the pre-trained object detector YOLOX[19].
+This integration involves using YOLOX generated proposal boxes as anchors to initialize object queries, a method that not only bolsters detection accuracy but also helps to alleviate conflicts between detection and association tasks, all the while maintaining the query propagation features inherent to MOTR.
+Similarly, MeMOT[10] innovates by incorporating track queries that model and predict target trajectories iteratively across video frames.
+This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
+Moreover, MOTRv3[9] and CO-MOT[20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework.
+MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance.
+In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces "shadows" to enhance the diversity and quantity of queries, further refining the MOT capabilities.
+2.2 Real Time Detection Method
+When discussing real-time detection models, most people initially think of YOLO[21], which transformed the problem of object detection into a regression model, achieving efficient object detection.
+The model divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions.
+The most advanced versions in the YOLO series currently are YOLOv8.2[22] and YOLOv9[23].
+Recently, the real-time detection model based on DETR, known as RT-DETR[24], has also demonstrated its extraordinary performance.
+RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the model structure.
+Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary.
+Therefore, inspired by the YOLO series[25, 22, 26], we designed a feature extraction network.
+3 DecoderTracker
+Figure 2: The overall structure of DecoderTracker.
 By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states.
 When the model is trained using the TALA strategy, the TBSP can be skipped.
-3.1 MO-YOLO Architecture
-MO-YOLO’s overall architecture is illustrated in the provided Fig.2.
-In this setup, a video sequence is input to YOLOv8’s backbone, neck to extract features.
-For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the RT-DETR decoder.
-The learning pattern of $q_{d}$ is similar to RT-DETR[21].
+Figure 3: The structure of IMFENet.
+It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences.
+For detailed comparisons, please refer to the appendix.
+3.1 DecoderTracker Architecture
+DecoderTracker’s overall architecture is illustrated in the provided Fig.2.
+In this setup, a video sequence is input to Image Feature Extraction Network(IMFENet) to extract features.
+For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
 For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder.
@@
 Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
+The structure of IMFENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction.
+The reason for not adopting the latest YOLOv9 is due to its slower training speed compared to YOLOv8.
+Through IMFENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
 3.2 Tracking Box Selection Process
-When training a detection model using the detection method and subsequently employing an end-to-end tracking approach, we observed a gradual increase in the number of tracking boxes for the same object over time.
-Upon examination, this phenomenon arises from the fact that each newly generated detect query is consistently assigned to an already tracked object.
-This is due to the nature of DETR, where a detect query can potentially be assigned to any object in the image, as label assignment is determined through binary matching of all detect queries and ground truth.
-In MOTR[42], a solution to this issue is introduced through the utilization of TALA.
-This strategy gradually suppresses the generation of detect queries similar to track queries during the training process, effectively addressing the identified problem.
-However, this process increases the number of parameters that the network needs to automatically adjust, resulting in a prolonged convergence time for the model.
-Therefore, this paper proposes a simple yet effective approach named tracking box selection process (TBSP).
-Experimental results demonstrate the efficacy of TBSP, revealing that this method allows the model to achieve comparable performance to fully trained models even without extensive training.
-Additionally, the simplicity of TBSP ensures that it introduces minimal additional computational burden.
-Figure 3: Refinement process for MO-YOLO on partially trained track queries.
-When MO-YOLO undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized.
-At that point, the track query generation method of MO-YOLO aligns with MOTR, as illustrated in the Fig.
-1.
+Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2
+The initial design of the TBSP was intended merely to equip a detection network with basic tracking functionality.
+We observed that when a detection network directly employs the MOTR[3] method to update track queries, the tracking boxes for the same object tend to increase over time.
+This accumulation was found to be due to each newly generated detect query invariably being assigned to an already tracked object.
+In the DETR, a detect query can be assigned to any object in the image, as label assignment is determined through bipartite matching between all detection queries and ground truth data.
+MOTR addresses this by utilizing tracklet-aware label assignment (TALA), which suppressively modulates the generation of detect queries similar to track queries during training, thus mitigating this issue.
+Consequently, we designed a simple filtering mechanism, TBSP, which enables the detection network to possess some tracking capabilities without additional training.
+However, given the network’s inherently weak tracking ability and the lack of robustness in TBSP, the tracking performance at this stage is quite limited.
+Nevertheless, we also discovered that using TBSP as a preprocessing step for TALA, and training the network in this fashion, can lead to some unexpected benefits.
+Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training.
+This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the model to optimize itself.
+For a discussion on why TBSP is effective, please refer to the appendix.
 We know that both track query and detect query contain information about the object’s bounding box(BBox).
@@
 Finally, the remaining detect queries are denoted as $\omega_{\det_{f}}^{t-1}$.
-Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2.
 Similarly, assuming the track query at frame $t$ (where $t>2$) is denoted as $\omega_{tr}^{t}$, it can be expressed as:
@@
 4.
-Incidentally, although TBSP makes MO-YOLO appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted.
-Simultaneously, to better illustrate the performance of our model compared to MOTR in CO-MOT[38] and MOTRv3[40], we have not adopted the labeling strategy used in CO-MOT and MOTRv3.
-3.3 Decoder Embeddings and Query Positions
-Differentiating itself from MOTR, MO-YOLO’s decoder employs a methodology akin to RT-DETR for its embeddings.
-In contrast to MOTR, MO-YOLO’s decoder input doesn’t originate from a distinct encoder; rather, it exploits RT-DETR’s IoU-aware query selection mechanism.
-Building upon prior works such as [35, 43, 50, 16], this process strategically selects high-quality image features as the initial object embeddings for the decoder, emphasizing IoU scores.
-It’s essential to clarify that in this context, the term "embeddings" specifically pertains to the initial embeddings of detection queries.
-The subsequent embeddings of track queries undergo generation through a TAN network, aligning with MOTR’s established process.
-To expedite model convergence, we implement a Query Positions generation strategy akin to CO-MOT.
-Drawing inspiration from DAB-DETR [19], we employ a fixed transformation to convert reference points into Query Positions.
-While experimentation with the RT-DETR approach involving a Multilayer Perceptron (MLP) to transform reference points shows no significant performance discrepancy, it is important to note that the use of MLP substantially increases training time.
-By the way, it is worth mentioning that this process differs significantly from MOTR.
-In MOTR, Query Positions are generated using a set of learnable parameters, whereas in MO-YOLO, they are generated using reference points.
+Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted.
+Simultaneously, to better illustrate the performance of our model compared to MOTR, we have not adopted the labeling strategy used in CO-MOT[8] and MOTRv3[9].
 3.3 Loss and Training Methods
-Due to the unique training mode adopted by MOTR, its batch size in a single GPU is constrained to 1.
-While this training approach enables the model to effectively learn the temporal information of the targets, it concurrently requires training the network to extract features of target appearances.
-To expedite the model training speed, a three-stage training strategy is employed in this study.
-In the first stage, the emphasis is placed on training the network to extract features of target appearances; the second stage focuses on learning temporal information; and the final stage comprehensively enhances the model’s performance by integrating the learning of temporal information and appearance features.
-This strategy is implemented to overcome the constraints posed by the unique training mode of MOTR, aiming to achieve more efficient model training.
+Due to MOTR’s similarity to RNN in its inference approach, its batch size is limited to 1.
+This often results in suboptimal utilization of GPU resources during training.
+Additionally, MOTR learns both image and motion features simultaneously.
+To expedite model training, this paper adopts a three-stage training strategy.
+In the first stage, emphasis is placed on training the network to extract target image features efficiently.
+The second stage involves training the model on weak tracking aspects using TBSP as a pretraining step for the third stage.
+Lastly, the third stage employs the TALA strategy to comprehensively enhance the model’s tracking performance.
 In the initial stage, the network is trained as a detection network, and the loss function is formulated defined as Eq.2:
 $\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)}{V_{i}}$ (2)
-$\mathcal{L}$ is the loss of single frame, which is similar to the detection loss in DETR, and it is shown in Eq.3.
-$\mathcal{L}\left(\left.\widehat{Y}_{t}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}$ (3)
-where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[17], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[26], $\lambda_{cls},\lambda_{l1},\lambda_{giou}$ are corresponding weight coefficients.
+$\mathcal{L}$ is the loss of single frame.
+Notably, during this stage, we employed the IoU-aware Query Selection from RT-DETR, so the loss function utilized is from RT-DETR.
+Unlike the original DETR, this stage introduces a loss termed $\mathcal{L}_{enc}$.
+This loss involves treating the initialized queries as boxes for the purpose of loss computation.
+The specifics are outlined below:
+$\mathcal{L}_{enc}(\hat{y},y)&=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},\hat{b},y,b)\\ &=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},c,IoU)\end{aligned}$ (3)
+where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively.
+Finally, the overall loss is as follows:
+$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (4)
+where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[27], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[28], $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients.
 In this stage, the network is capable of initially learning the appearance features of objects to be tracked.
@@
 The loss functions used in this stage and the third stage are as Eq.5.
-Simultaneously, it is important to note that, at this juncture, the determination of the newly generated track queries relies solely on the scores of object queries and whether they have been filtered out by TBSP, as more advanced strategies like TALA have not been employed.
 $\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\sum_{n=1}^{N}{\left(\mathcal{L}\left(\left.\widehat{Y}_{tr}^{t}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)+\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)\right)}}{\sum_{n=1}^{N}{\left(V_{t}\right)}},N=5$ (5)
@@
 $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively.
-The third stage of training involves the utilization of TALA policy during the training phase, while simultaneously deactivating the TBSP.
+At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects.
+Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage.
+The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP.
 Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance.
-This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, the TBSP policy can be turned off during inference with negligible impact on overall performance.
-In addition, it is important to note that while the first and second stages of MO-YOLO share similarities in generating a new track query from the object query, discrepancies may arise in the third stage.
-Tracking objects with the track query does not involve additional techniques, and the process across the three stages of MO-YOLO closely resembles that of MOTR.
+This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
 4 Experiments
 4.1 Datasets and Metrics
-This paper introduces MO-YOLO, an innovative MOT method proficient in addressing intricate scenarios involving occlusions, interactions, and diverse motions.
-To validate its performance, we evaluate the algorithm on three demanding datasets: DanceTrack[29], MOT17[24], and KITTI[12].
+To validate DecoderTracker’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[32], MOT17[33], and KITTI[34].
 DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures.
@@
 4.2 Implementation Details
-The rationale for adopting MO-YOLO originates from our observation of the high training costs and suboptimal inference speed associated with MOTR.
+The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR.
 Thus, our aim is to propose a new baseline model surpassing MOTR’s performance.
 To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [10], MOTRv2 [18], MOTRv3 [9], and CO-MOT [8].
-To achieve a comparable parameter count with MOTR, we simultaneously utilize YOLOv8-L [14] and RT-DETR-L [21] in constructing MO-YOLO (henceforth referred to as MO-YOLO-L concerning model size).
-The initial weights are obtained from a model pre-trained on the COCO dataset [18].
 The majority of experiments were conducted using PyTorch on a single NVIDIA GeForce RTX 4090 GPU.
@@
 Subsequent stages utilized a uniform batch size of 1.
-For MOT17, like MOTR, we integrated extra data from the CrowdedHuman dataset [28], employing random displacement to generate video segments with pseudo-trajectories.
+For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [37], employing random displacement to generate video segments with pseudo-trajectories.
 Training comprised 120 epochs in the first stage, followed by 30 epochs in the second.
 The third stage extended to 20 epochs with TBSP, or 55 without.
-For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 6, and the third, if TBSP was used, was 6 epochs, or 15 without TBSP.
+For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP.
 For KITTI, initial training involved 80 epochs, followed by 25 in the second stage.
@@
 Additionally, approximately 5k images from BDD100k [38] were selected for additional training.
-Note: With the exception of Table 4(b), all data concerning MO-YOLO in this paper pertains to MO-YOLO-L.
-Additionally, unless stated otherwise, the provided data pertains to the final end-to-end model.
 4.3 Performance Evaluation and Comparison
-In Table 1, a comprehensive comparison of training time and speed was conducted between MO-YOLO and the MOTR series.
+In Table 1, a comprehensive comparison of training time and speed was conducted between DecoderTracker and the MOTR series.
+It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations.
+This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
 (a) Comparison of Training Time on MOT17
@@
 MOTR[3] 8 2080ti $\sim$ 96.0 hours
-MO-YOLO stage1 1 2080ti 12.1 hours
-MO-YOLO stage2 1 2080ti 30.8 hours
-MO-YOLO(TBSP) 1 2080ti 43.4 hours
-MO-YOLO 1 2080ti 68.7 hours
+DecoderTracker stage1 1 2080ti 12.1 hours
+DecoderTracker stage2 1 2080ti 30.8 hours
+DecoderTracker(TBSP) 1 2080ti 43.4 hours
+DecoderTracker 1 2080ti 68.7 hours
 (b) Comparison of Speed on Dancetrack
@@
 MOTRv3[9] 10.6
-MO-YOLO(TBSP) 18.5
-MO-YOLO 19.6
-Table 1: Comparison of training time (including all previous stages) and speed between MO-YOLO and MOTR series on a single V100 GPU.
+DecoderTracker(TBSP) 18.5
+DecoderTracker 19.6
+Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU.
 The data for the MOTR series is sourced from [3], [9] or their open source address.
-It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations.
-This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
-MO-YOLO demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed.
-The primary reason behind this efficiency gain lies in the architectural simplicity of MO-YOLO.
-Unlike MOTR, MO-YOLO adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design.
-This reduction in complexity enables MO-YOLO to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU.
+DecoderTracker demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed.
+The primary reason behind this efficiency gain lies in the architectural simplicity of DecoderTracker.
+Unlike MOTR, DecoderTracker adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design.
+This reduction in complexity enables DecoderTracker to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU.
 In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training.
-The absence of an encoder in MO-YOLO contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
-On the Dancetrack test dataset, MO-YOLO(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3).
-Particularly noteworthy is the remarkable 19.6 FPS achieved by MO-YOLO, emphasizing its outstanding performance in inference speed.
-Table 2: Performance comparison between MO-YOLO and existing methods on the Dancetrack[29] dataset under the private detection protocols.
-Numbers are highlighted in bold when they correspond to either MO-YOLO or the superior metric from MOTR.
-Methods HOTA AssA DetA IDF1 MOTA
+The absence of an encoder in DecoderTracker contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
+Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[32] benchmark.
+Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR.
+Methods HOTA IDF1 DetA AssA MOTA
 Non-End-to-end:
-ByteTrack[45] 47.3 31.3 71.6 52.5 89.5
-GTR[49] 48.0 31.9 72.5 50.3 84.7
-DST-track[6] 51.9 34.6 82.3 51.0 84.9
-OC-SORT[5] 54.6 38.0 80.4 54.2 89.4
+ByteTrack[14] 47.3 52.5 71.6 31.3 89.5
+GTR[39] 48.0 50.3 72.5 31.9 84.7
+DST-track[40] 51.9 51.0 82.3 34.6 84.9
+OC-SORT[41] 54.1 54.2 80.4 38.0 89.4
 End-to-end:
-MeMOTR[10] 63.4 52.3 77.0 65.5 85.4
-MOTR[42] 54.2 40.2 73.5 51.5 79.7
-MO-YOLO(Ours) 52.0 35.7 75.9 51.4 83.5
-Table 2, Table 3, and Table 4 exhibit the final results of end-to-end MO-YOLO on the Dancetrack, MOT17, and KITTI datasets, respectively.
-In contrast to MOTR, MO-YOLO demonstrates competitive performance across multiple metrics on both datasets.
+MeMOTR[10] 63.4 65.5 77.0 52.3 85.4
+MOTR[3] 54.2 51.5 73.5 40.2 87.6
+DecoderTracker(Ours) 54.2 54.2 78.7 37.4 87.8
+On the Dancetrack test benchmark, DecoderTracker(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3).
+Particularly noteworthy is the remarkable 19.6 FPS achieved by DecoderTracker, emphasizing its outstanding performance in inference speed.
+Table 2, Table 3, and Table 4 exhibit the final results of end-to-end DecoderTracker on the Dancetrack, MOT17, and KITTI benchmark, respectively.
+In contrast to MOTR, DecoderTracker demonstrates competitive performance across multiple metrics on both datasets.
 It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3.
-The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our MO-YOLO.
+The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our DecoderTracker.
 MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training.
-Consequently, comparing its tracking performance directly with that of MO-YOLO and MOTR may not be entirely fair.
-Table 3: Performance comparison between MO-YOLO and existing methods on the MOT17[24] dataset.
-method HOTA AssA DetA IDF1 MOTA
+Consequently, comparing its tracking performance directly with that of DecoderTracker and MOTR may not be entirely fair.
+Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17[33] set.
+method HOTA IDF1 DetA AssA MOTA
 Non-End-to-end:
-FairMOT[46] 59.3 58.0 60.9 72.3 73.7
-ByteTrack[45] 62.8 62.2 63.8 77.2 78.9
-StrongSORT[49] 63.5 63.7 63.6 78.3 78.5
-UTM[39] 64.0 62.5 65.9 78.7 81.8
-OC-SORT[5] 63.2 63.2 63.2 77.5 78.0
-Deep OC-SORT[22] 64.9 65.9 64.1 80.6 79.4
+FairMOT[42] 59.3 72.3 60.9 58.0 73.7
+ByteTrack[14] 62.8 77.2 63.8 62.2 78.9
+StrongSORT[39] 63.5 78.3 63.6 63.7 78.5
+UTM[43] 64.0 78.7 65.9 62.5 81.8
+OC-SORT[41] 63.2 77.5 63.2 63.2 78.0
+Deep OC-SORT[6] 64.9 80.6 64.1 65.9 79.4
 End-to-end:
-MeMOTR[10] 58.8 58.4 59.6 69.0 72.5
-MOTR[42] 57.2 55.8 58.9 59.4 71.9
-MO-YOLO(Ours) 55.5 55.9 55.8 70.1 66.8
-For Dancetrack (Table 2), MO-YOLO achieves comparable results to MOTR, with slightly lower scores in HOTA and AssA but outperforming MOTR in DetA.
-Specifically, MO-YOLO achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
-In the MOT17 (Table 3), MO-YOLO demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA.
-Despite trailing slightly behind MOTR in these metrics, MO-YOLO compensates with a comparable DetA and excels in IDF1 by almost 10 points.
-This advantage can be attributed to MO-YOLO’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand.
-Additionally, MO-YOLO maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
-Table 4: Performance comparison between MO-YOLO and existing methods on the KITTI[12] dataset.
-Car Pedestrian
+MeMOTR[10] 58.8 69.0 59.6 58.4 72.5
+MOTR[3] 57.2 59.4 58.9 55.8 71.9
+DecoderTracker(Ours) 55.7 70.5 55.7 56.0 66.5
+For Dancetrack (Table 2), DecoderTracker achieves comparable results to MOTR, with same scores in HOTA but outperforming MOTR in DetA.
+Specifically, DecoderTracker achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
+In the MOT17 (Table 3), DecoderTracker demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA.
+Despite trailing slightly behind MOTR in these metrics, DecoderTracker compensates with a comparable DetA and excels in IDF1 by almost 10 points.
+This advantage can be attributed to DecoderTracker’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand.
+Additionally, DecoderTracker maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
+Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI[34] set.
+Pedestrian Car
 Tracker HOTA MOTA AssA HOTA MOTA AssA
 Non-End-to-end:
-CenterTr[48] 73.0 88.8 71.1 40.3 53.8 36.9
-PermaTr[30] 77.4 90.8 77.6 47.4 65.0 43.6
-PolarMOT[15] 75.1 80.0 76.9 43.6 46.9 48.1
-TripletTrack[23] 73.6 84.3 74.7 50.0 46.5 57.8
+CenterTr[44] 40.3 53.8 36.9 73.0 88.8 71.1
+PermaTr[45] 47.4 65.0 43.6 77.4 90.8 77.6
+PolarMOT[46] 43.6 46.9 48.1 75.1 80.0 76.9
+TripletTrack[47] 50.0 46.5 57.8 73.6 84.3 74.7
 End-to-end:
-MO-YOLO 72.2 83.2 73.8 51.5 56.8 58.4
-Table 4 provides an overview of MO-YOLO’s performance on the KITTI dataset.
-MO-YOLO achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively.
-These results position MO-YOLO as a strong contender in multi-object tracking on the KITTI.
-In conclusion, MO-YOLO stands out as a highly advantageous model in the realm of multi-object tracking, demonstrating superior object detection accuracy and overall tracking performance compared to MOTR, as evidenced by competitive results on diverse datasets, including Dancetrack, MOT17, and KITTI.
-The notable improvements observed in MeMOTR relative to MOTR suggest that MO-YOLO has significant untapped potential for performance enhancement.
-Positioned as a robust alternative, MO-YOLO can serve as an excellent new baseline, driving advancements in end-to-end models within the field of multi-object tracking.
+DecoderTracker 51.5 56.8 58.4 72.2 83.2 73.8
+Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark.
+DecoderTracker achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively.
+These results position DecoderTracker as a strong contender in MOT on the KITTI.
 4.4 Ablation Study
-To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack dataset.
+To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark.
 The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set.
-Additionally, the impact of two methods for generating Query Positions—utilizing the formula from DAB-DETR[19] or employing an MLP—on the model’s performance was investigated.
-Notably, MO-YOLO-L, MO-YOLO-M, and MO-YOLO-S differ based on their foundations in YOLOv8-L,YOLOv8-M, and YOLOv8-S, respectively, providing insights into model variations and their implications.
-(a) Performance of MO-YOLO at different training stages Training Stage HOTA AssA MOTA Initial 38.5 21.6 76.8 Intermediate 44.5 25.5 84.2 Final (TBSP) 51.9 34.6 84.9 Final 53.6 35.9 84.7 (b) Performance across different model sizes Model HOTA MOTA FPS MO-YOLO-L 53.6 84.7 19.6 MO-YOLO-M 49.7 82.9 22.7 MO-YOLO-S 42.9 79.3 26.2
-(c) The impact of $\delta_{iou}$ on model performance when using TBSP $\delta_{iou}$ HOTA AssA MOTA 0.3 49.7 32.3 79.1 0.5 51.9 34.6 84.9 0.7 51.3 34.8 83.2 0.9 50.0 32.8 81.3 (d) Effect of query position generation strategy Strategy HOTA MOTA FPS MLP 53.3 85.1 18.5 Formula 53.6 84.7 19.6
-Table 5: Ablation Studies on Our Proposed MO-YOLO on the DanceTrack validation set.
-The performance of MO-YOLO improves consistently across different training stages, peaking in the final stage.
-Different model sizes (MO-YOLO-L, MO-YOLO-M, MO-YOLO-S) achieve a balance between accuracy and computational efficiency.
-The similarity in their speeds is likely attributable to the consistent scale of the decoder.
+Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR[48] or employing an MLP, on the model’s performance was investigated.
+(a) Performance of DecoderTracker at different training stages
+Training Stage HOTA AssA MOTA
+First 38.5 21.6 76.8
+Second 44.5 25.5 84.2
+Final (TBSP) 51.9 34.6 84.9
+Final 53.6 35.9 84.7
+(b) The presence or absence of TAN during the second training stages can significantly impact the performance of the model.
+Model HOTA AssA MOTA
+Second 44.5 25.5 84.2
+Second(w/o) 42.1 21.4 80.8
+Final 53.6 35.9 84.7
+Final(w/o) 53.4 36.1 84.5
+(c) The impact of $\delta_{iou}$ on model performance when using TBSP
+$\delta_{iou}$ HOTA AssA MOTA
+0.3 49.7 32.3 79.1
+0.5 51.9 34.6 84.9
+0.7 51.3 34.8 83.2
+0.9 50.0 32.8 81.3
+(d) Effect of query position generation strategy
+Setting HOTA MOTA FPS
+sin/cos 53.3 85.1 18.5
+MLP 53.6 84.7 19.6
+Table 5: Ablation Studies on Our Proposed DecoderTracker on the DanceTrack validation set.
+The performance of DecoderTracker improves consistently across different training stages, peaking in the final stage.
 The choice of $\delta_{iou}$ impacts precision and recall trade-offs, with higher values leading to a slight decrease in performance.
-Position generation strategies, MLP and Formula, yield similar performance in terms of detection accuracy, with slight differences in FPS.
-Overall, the ablation studies demonstrate the effectiveness of MO-YOLO across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
-Figure 5: MO-YOLO and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 dataset.
+Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS.
+Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
+It is noteworthy that in Table 5(b), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages.
+It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance.
+However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
 5 Conclusion and Limitation
-This paper presents MO-YOLO, an efficient and lightweight end-to-end multi-object tracking model that optimizes computational resources.
-By integrating YOLO and RT-DETR principles, the study establishes a novel tracking paradigm with accelerated training strategies.
-While MO-YOLO outperforms MOTR in resource efficiency, it marginally lags behind in performance, attributed to factors like YOLO’s detection capabilities and reduced input image dimensions ($640\times 640$).
-Despite enhancing computational efficiency, the dimension reduction raises concerns about potential information loss, impacting overall effectiveness.
-Additionally, latent issues from MOTR may surface in MO-YOLO, as depicted in Fig.5.
-This research contributes to advancing real-time computer vision applications.
-Appendix 0.A Appendix
-0.A.1 Disclaimer
-The data used in this paper is sourced from publicly available datasets related to human subjects.
-While every effort has been made to ensure the accuracy and reliability of the data, the authors make no representations or warranties of any kind, express or implied, about the completeness, accuracy, reliability, suitability, or availability with respect to the data.
-The authors shall not be liable for any loss or damage, including but not limited to indirect or consequential loss or damage, arising from the use of the data in this research paper.
-Users of the data are encouraged to independently verify the accuracy and reliability of the data before making any decisions or taking any actions based on it.
-0.A.2 More Comprehensive Analysis of Experimental Results and Model Parameter
-The initial motivation for building MO-YOLO was not because of its slow inference speed, but rather because of its high training resource requirements.
-With 8 V100 GPUs using pre-trained weights, it took approximately 84 hours to train 200 epochs on an extended MOT17[24] with only about 10,000 images.
-Each epoch took around 0.42 hours, and on average, each GPU could train on approximately 3,000 images per hour.
-It is worth noting that MOTR[42] is based on Deformable-DETR[50], which was trained on the MSCOCO dataset[18] with approximately 110,000 images.
-In single-scale mode (which can be considered similar to MOTR’s training), using eight V100 GPUs, each epoch took approximately 0.4 hours, and on average, each GPU could train on around 30,000 images per hour.
-Upon careful consideration of the reasons, the first aspect to be contemplated is the batch size.
-The collective average loss (CAL) mechanism and the tracklet-aware label assignment (TALA) strategy employed in MOTR enable the network to learn some latent temporal information, consequently restricting the batch size on a single card to 1.
-Setting the batch size to 1 may result in underutilization of the corresponding computational resources (e.g., GPU, TPU, or other possible devices).
-Additionally, MOTR is trained in a one-shot manner, requiring the model not only to grasp motion information but also to learn object appearance features.
-Hence, we proposed a multi-step training strategy to first allow the model to learn appearance features before proceeding to learn object motion characteristics.
-Experimental results confirm the effectiveness of this strategy.
-As for the TBSP, its initial purpose was simply to provide the detection network with a basic tracking capability.
-However, it also pleasantly surprised us by further reducing training time.
-Here, we also provide the required rounds and results if TALA were to be used directly without the TBSP strategy, as shown in Table 6.
-The training here is conducted on a single NVIDIA 4090 GPU.
-And this training process is not included in the first stage
-Table 6: Comparison of MO-YOLO with and without using TBSP on the Dancetrack val dataset.
-Initial Intermediate Final HOTA MOTA Epochs Total Time
-53.6 84.7 27 56.7 hours
-52.9 83.8 40 85.2 hours
-In order to provide a more comprehensive illustration of the characteristics of our model, we list the sizes of different models, namely MO-YOLO-L (45.8M), MO-YOLO-M (29.9M), and MO-YOLO-S (19.7M), as depicted in Fig.
-6.
-Figure 6: The model sizes of MOTR and MO-YOLO, as well as their FPS and HOTA on the Dancetrack.
-The table compares the performance of MO-YOLO with and without using the TBSP strategy on the Dancetrack validation dataset across different training phases: Initial, Intermediate, and Final.
-Comparison of the two rows reveals that employing TBSP throughout all phases leads to slightly better performance in terms of HOTA and MOTA, with a significantly lower total training time.
-This underscores the importance of the TBSP strategy in reducing the overall training duration while maintaining or even improving tracking accuracy.
-We hypothesize that the efficacy of the TBSP strategy may stem from its endowment of the model with a degree of autonomy, facilitating the preliminary acquisition of object motion information that would otherwise necessitate numerous training iterations.
-Analogously, the strategy resembles the use of training wheels in bicycle riding, which assist novices in achieving smooth operation while providing an opportunity to experience the sensation of riding.
-Consequently, this approach aids in the comprehensive mastery of the skill.
-We posit that a comparable scenario may manifest within this context.
-0.A.3 RT-MOTR
-At the same time, to demonstrate the advantages of our MO-YOLO models constructed using YOLO and RT-DETR relative to end-to-end tracking models built directly on RT-DETR, we have also constructed such an end-to-end tracking model based on RT-DETR (referred to as RT-MOTR), and conducted performance testing on it.
-NOTE: in Table 7, both the train time and epochs refer to the second and third stages like Table 6.
-For the first stage, both models were trained for 15 epochs, and the training device is a single 4090, while the testing device is a single V100.
-The table compares the performance of MO-YOLO and RT-MOTR on the Dancetrack validation dataset.
-It presents various metrics including HOTA (Higher Order Tracking Accuracy), training epochs, training speed (hours per epoch), total training time, and frames per second (FPS).
-MO-YOLO achieves a HOTA of 53.6 with 27 training epochs, an average training speed of 2.08 hours per epoch, a total training time of 56.7 hours, and an FPS of 19.6.
-In comparison, RT-MOTR achieves a HOTA of 53.4 with 35 training epochs, an average training speed of 2.27 hours per epoch, a total training time of 90.8 hours, and an FPS of 20.3.
-Comparing the performance metrics of the two methods, we observe that MO-YOLO slightly outperforms RT-MOTR in terms of HOTA, while also having a shorter training time albeit with a slightly lower FPS.
-This indicates that MO-YOLO exhibits advantages in both tracking accuracy and training efficiency.
-Table 7: Comparison of MO-YOLO and RT-MOTR on the Dancetrack val dataset.
-Methods HOTA Train Epochs Speed (hours/epoch) Train Times FPS
-MO-YOLO 53.6 27 2.08 56.7 h 19.6
-RT-MOTR 53.4 35 2.27 79.5 h 20.3
-Regarding why the speeds of these two models are comparable, our experimental findings suggest that the primary time-consuming component is the decoder responsible for tracking.
-Therefore, optimizing the decoder represents a crucial area for future investigation.
-This outcome underscores the rationale behind integrating elements from YOLOv8 and RT-DETR to construct a rapid end-to-end tracking network.
-Furthermore, the observed disparity in training efficiency between MO-YOLO and RT-DETR can be attributed to the former’s notably simpler architecture.
-Moreover, our decision to adopt YOLO as the backbone of the network is motivated by its propensity for swifter updates compared to DETR-based models.
-Recent developments such as YOLOv9[34] exemplify this trend.
-While YOLOv9 may trade off training time for enhanced precision and speed (resulting in at least a twofold increase in training duration compared to YOLOv7[33]), we anticipate future iterations will offer refined improvements.
-Such advancements will, in turn, facilitate further enhancements to MO-YOLO.
+This paper presents an efficient decoder-only end-to-end MOT (Multi-Object Tracking) model that boasts faster inference speeds and reduced training times.
+Additionally, this paper introduces new training strategies aimed at accelerating model convergence.
+Although device limitations prevented us from validating our training strategies on the original MOTR model, we believe these strategies are effective.
+This research provides a more efficient benchmark model for the multi-object tracking field compared to MOTR, laying a foundation for the emergence of real-time end-to-end MOT models and offering insights for future development.
+In subsequent work, we should further explore more efficient network architectures to construct real-time MOT models.
+Of course, we also plan to incorporate techniques from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to enhance tracking performance.
+Appendix A More Discusstion About TBSP
+Figure 5: Refinement process for DecoderTracker on partially trained track queries.
+When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized.
+At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Fig.
+1.
+Table 6: Comparison of DecoderTracker with and without using TBSP on the Dancetrack val set.
+Setting HOTA MOTA Epochs Total Time
+TBSP 53.6 84.7 27 56.7 hours
+w/o 53.4 84.2 40 85.2 hours
+The initial reason for developing the MO-YOLO model was not due to slow inference speed, but rather the extensive training resources required.
+Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs.
+Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
+Fortunately, our DecoderTracker model has addressed this issue.
+Although its inference speed is not yet real-time, its convergence rate is already satisfactory.
+Although we have not achieved real-time inference speeds, we have significantly improved the speed of MOTR.
+Regarding the TBSP, its original purpose was simply to endow the detection network with basic tracking functionality.
+However, it also provided some unexpected benefits, such as further reducing training time.
+Below, we list the results and required iterations if training was conducted directly using TALA, as shown in Table 6.
+We believe the effectiveness of the TBSP strategy may be due to its function as a form of weak or self-supervised training for motion features, compared to TALA.
+Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence.
+For instance, in the NLP field, large language models like BERT[49] and GPT[1] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns.
+Subsequently, these models undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering.
+This approach has been empirically proven to significantly boost the performance of LLMs.
+In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[50, 51].
+Typically, these models are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
+We believe that the TBSP, a weak supervision process, enables the model to preliminarily learn object motion information that would otherwise require many more training iterations.
+It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill.
+A similar situation may exist here.
+Although TBSP did not enhance model performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
+Appendix B Design of IMFENet
+The design of this feature network is intended to enhance the model’s efficiency and increase its inference speed.
+Thus, our design primarily references the network structure of YOLOv8.
+The backbone structure of our model is almost identical to that of YOLOv8.
+Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
+Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M.
+This change is specifically reflected in the BottleNeckF module.
+In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load.
+These modifications are illustrated in Figure 6, and their impact on performance is shown in Table 7.
+Figure 6: The difference between our proposed model and YOLOv8 in terms of modules. Our modules are more lightweight. Specific model details can refer to the figure.3.
+Figure 7: DecoderTracker and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 benchmark.
+Table 7: The impact of our network and YOLOv8 structure on DecoderTracker performance on the Dancetrack val set. Training time does not include the first stage
+Setting HOTA IDF1 MOTA AssA FPS
+IMFENet 53.6 53.8 84.7 35.9 19.6
+YOLOv8 52.4 51.4 84.2 36.1 18.9
+Appendix C Limitation
+Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection models.
+Both models fail under similar circumstances, as illustrated in Figure.7.
+This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the model.
+Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR.

### v5

@@
-DecoderTracker: Decoder-Only Method for Multiple-Object Tracking
+DecoderTracker: Decoder-Only End-To-End method for Multiple-Object Tracking
 Abstract
-Decoder-only models, such as GPT[1], have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer models.
-Over the years, end-to-end models based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking.
-However, the significant computational resource consumption of these models leads to less friendly inference speeds and training times.
-To address these issues, this paper attempts to construct a lightweight Decoder-only model: DecoderTracker for end-to-end multi-object tracking.
-Specifically, drawing on some real-time detection models, we have developed an image feature extraction network which can efficiently extract features from images to replace the encoder structure.
-In addition to minor innovations in the network, we analyze the potential reasons for the slow training of MOTR-like models and propose an effective training strategy to mitigate the issue of prolonged training times.
-On the DanceTrack dataset, without any bells and whistles, DecoderTracker’s tracking performance slightly surpasses that of MOTR, with approximately twice the inference speed.
-Furthermore, DecoderTracker requires significantly less training time compared to MOTR.
+Decoder-only methods, such as GPT, have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer methods.
+Over the years, end-to-end methods based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking.
+However,The substantial computational resource consumption of these methods, coupled with the optimization challenges posed by dynamic data, results in less favorable inference speeds and training times.
+To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTrack, a novel end-to-end tracking method.
+Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTrack+ by incorporating a Fixed-Size Query Memory and refining certain attention layers.
+Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, with inference speeds 2.06 and 3.03 times faster than MOTR, respectively.
+keywords:
+Multi-object tracking , Decoder , End-to-end
+journal: Pattern Recognition11footnotetext: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
 1 Introduction
 Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images.
-The success of transformer[2] models in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT models like MOTR[3] have been introduced in the MOT domain.
-These models offer advantages over the widely used tracking-by-detection methods[4, 5, 6, 7], such as eliminating the need for manual design or selection of features.
-Additionally, some enhanced methods[8, 9, 10] of MOTR have demonstrated the potential of this model class in terms of tracking accuracy.
-However, the inference speed of these models is often suboptimal, and their RNN[11] like linear inference structure makes them less GPU friendly during training.
+The success of transformer [1] methods in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT methods like MOTR [2] have been introduced in the MOT domain.
+These methods offer advantages over the widely used tracking-by-detection methods [3, 4, 5, 6], such as eliminating the need for manual design or selection of features.
+Additionally, some enhanced methods [7, 8, 9] of MOTR have demonstrated the potential of this method class in terms of tracking accuracy.
 […]
-At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the model to engage in less stringent learning might also be effective.
+Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios.
+The number of detect queries is typically fixed, while the number of track queries fluctuates throughout the tracking process.
+This discrepancy results in dynamic data during inference, which introduces optimization challenges that limit the actual runtime speed of the method.
+To address this, we propose the Fixed-Size Query Memory (FSQM), which pads the tracking query sequence with empty queries to maintain a consistent number of queries during inference.
+By eliminating dynamic data, we have developed DecoderTracker+, which has significantly improved inference speed compared to MOTR and DecoderTracker.
+At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the method to engage in less stringent learning might also be effective.
 Is there a possibility that this process could yield unexpected benefits?
-Based on this idea, we proposed the Tracking Box Selection Process (TBSP).
-Although the methods we propose have not yet enabled the inference speed of MOTR-like models to achieve real-time performance, we believe that this paper could provide valuable insights for future researchers.
+It is important to note that weak supervision, or self-supervised learning, has achieved significant success in many computer vision tasks [12, 13].
+Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during the training.
+Thus, DecoderTracker has been constructed by us.
+After completing the aforementioned architectural optimizations mentioned above, we observed that while the training duration was significantly optimized, the inference speed did not improve as expected.
+This phenomenon prompted us to delve deeper into the issue.
+In theory, streamlining the network architecture should more notably enhance inference speed rather than training efficiency.
+To investigate this anomaly, we systematically analyzed the time consumption distribution across various modules.
+The results revealed that the decoder module accounted for the majority of the time, consuming nearly two-thirds of the model’s total inference time, which is evidently unreasonable.
+Through further cross-model comparative analysis, we identified that this issue also exists in MOTR but is absent in its base model, Deformable DETR.
+Controlled variable experiments indicated that fixing the number of queries significantly reduces decoder latency.
+Considering factors such as GPU memory usage, we ultimately pinpointed the root cause to be the dynamic query mechanism.
+This mechanism introduces unnecessary overhead during inference, such as inefficient GPU memory allocation in modern deep learning frameworks ,e.g., PyTorch/TensorFlow, and may also hinder compilation optimizations like operator fusion due to dynamic computation graphs.
+To address this issue, we proposed the Fixed-Size Query Memory (FSQM) for handling fixed-query tracking.
+By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTrack+.
+Experimental results demonstrated that, although DecoderTrack+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
 In summary, the contributions of this paper are as follows:
@@
   • 2)
-    A second key innovation introduces a unique training strategy characterized by a three-stage process.
-This strategy accelerates model convergence, reducing overall training time and enhancing training process efficiency.
+    The second key innovation involves the introduction of a unique training strategy characterized by a three-stage process.
+At the same time, we incorporated TBSP (Temporal Balanced Sampling and Propagation), a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods.
+These strategies accelerate method convergence, reduce overall training time, and enhance the efficiency of the training process.
   • 3)
-    The third contribution introduces TBSP, a strategy that can serve as a preliminary weakly supervised training approach for end-to-end MOT model training.
+    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTrack+, which addresses the additional latency caused by dynamic data.
+This provides a fully viable technical pathway for the engineering deployment of MOTR-like end-to-end MOT models.
 2 Related work
-Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios.
-The track query set, initially empty, is dynamically adjusted in length.
-Detect queries play a role in identifying new objects.
 2.1 MOTR Series
-Unlike tracking-by-detection models exemplified by SORT[4, 14, 5], and previous transformer-based MOT models[15, 16], MOTR[3] is the first model to achieve end-to-end multi-object tracking.
+Unlike tracking-by-detection methods exemplified by SORT [3, 14, 4], and previous transformer-based MOT methods [15, 16], MOTR [2] is the first method to achieve end-to-end multi-object tracking.
 It extends the object query from DETR [17] into a track query, thereby enabling tracking capabilities.
 It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing.
-Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for modeling long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration.
+Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for methoding long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration.
 These features establish MOTR as a significant advancement in multi-object tracking technology.
@@
 This integration involves using YOLOX generated proposal boxes as anchors to initialize object queries, a method that not only bolsters detection accuracy but also helps to alleviate conflicts between detection and association tasks, all the while maintaining the query propagation features inherent to MOTR.
-Similarly, MeMOT[10] innovates by incorporating track queries that model and predict target trajectories iteratively across video frames.
+Similarly, MeMOT [9] innovates by incorporating track queries that method and predict target trajectories iteratively across video frames.
 This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
-Moreover, MOTRv3[9] and CO-MOT[20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework.
+MOTRv3 [8] and CO-MOT [20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework.
 MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance.
-In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces "shadows" to enhance the diversity and quantity of queries, further refining the MOT capabilities.
+In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces ”shadows” to enhance the diversity and quantity of queries, further refining the MOT capabilities.
+2.2 Tracking by Detection
+Tracking by Detection are among the most classical and widely used types of MOT. These methods typically employ a detector to first identify objects and then use various other methods for tracking.
+In past years, the main methods within the Tracking by Detection category were SORT [3] and its derivative methods.
+The primary approach of these methods is to use a Kalman Filter to predict the positions of targets in the next frame and then perform data association with the detection results provided by the detector.
+Many subsequent improvements to these methods [5, 14, 4, 21] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association.
+There are also a series of methods like OC-SORT [22] that purely improve the Kalman Filter-based tracking method using kinematic methods.
+In recent years, methods such as SUSHI [23], Glan [24] and PuTR [25] have emerged, which employ networks for tracking post-detection.
+The former two utilize Graph Neural Networks (GNN), while the last is based on transformer.
 2.3 Real Time Detection Method
-When discussing real-time detection models, most people initially think of YOLO[21], which transformed the problem of object detection into a regression model, achieving efficient object detection.
-The model divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions.
-The most advanced versions in the YOLO series currently are YOLOv8.2[22] and YOLOv9[23].
-Recently, the real-time detection model based on DETR, known as RT-DETR[24], has also demonstrated its extraordinary performance.
-RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the model structure.
-Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary.
-Therefore, inspired by the YOLO series[25, 22, 26], we designed a feature extraction network.
-3 DecoderTracker
 Figure 2: The overall structure of DecoderTracker.
 By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states.
-When the model is trained using the TALA strategy, the TBSP can be skipped.
-Figure 3: The structure of IMFENet.
+To put it simply, $q_{tr}$ represents the objects in the video sequence that need to be tracked continuously, while $q_{d}$ represents the new objects appearing in the video frame.
+When the method is trained using the TALA strategy, the TBSP can be skipped.
+The process of DecoderTracker+ is consistent with this scenario, but the main difference lies in the updating of $q_{tr}$ and whether it is of fixed length.
+The FSQM of DecoderTracker+ pads $q_{tr}$ with all-zero queries.
+This approach effectively addresses optimization issues caused by dynamic data.
+When discussing real-time detection methods, most people initially think of YOLO [26], which transformed the problem of object detection into a regression method, achieving efficient object detection.
+The method divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions.
+The most advanced versions in the YOLO series currently are YOLOv8.2 [27] and YOLOv9 [28].
+Recently, the real-time detection method based on DETR, known as RT-DETR [29], has also demonstrated its extraordinary performance.
+RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the method structure.
+Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary.
+Therefore, inspired by the YOLO series [28, 27], we designed a feature extraction network.
+3 Method
+3.1 DecoderTracker
+Figure 3: The structure of FENet.
 It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences.
-For detailed comparisons, please refer to the appendix.
-3.1 DecoderTracker Architecture
-DecoderTracker’s overall architecture is illustrated in the provided Fig.2.
-In this setup, a video sequence is input to Image Feature Extraction Network(IMFENet) to extract features.
+Specific module details can be found in Figure 4.
+The extracted features will be fed into the Decoder and introduced to the cross-attention module to interact with the queries.
+Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder and the TAN.
+3.1.1 DecoderTracker Architecture
+DecoderTracker’s overall architecture is illustrated in the provided Figure 2.
+In this setup, a video sequence is input to image Feature Extraction Network(FENet) to extract features.
 For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
@@
 Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
-The structure of IMFENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction.
-The reason for not adopting the latest YOLOv9 is due to its slower training speed compared to YOLOv8.
-Through IMFENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
-3.2 Tracking Box Selection Process
-Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2
+3.1.2 Design of FENet
+The structure of FENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction.
+The reason for not adopting the latest YOLOv9 or YOLOv10 is due to its slower training speed compared to YOLOv8.
+Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
+The design of this FENet is intended to enhance the method’s efficiency and increase its inference speed.
+Thus, our design primarily references the network structure of YOLOv8.
+The backbone structure of our method is almost identical to that of YOLOv8.
+Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
+Figure 4: Specific module details, modules not specified in the diagram are network structures or APIs that come with PyTorch.
+These modules are primarily constructed based on CNN, capable of efficiently extracting image features.
+The employment of residual connections ensures that issues like gradient explosion and vanishing gradients are largely avoided.
+3.1.3 Tracking Box Selection Process
+Algorithm 1 TBSP flow
+0:  $\hat{\omega}_{tr}^{t-1}$: Track queries assignment, $\hat{\omega}_{det}^{t-1}$: Detect queries assignment, $\delta_{iou}$: IOU threshold
+1:  $[BBox_{1}^{tr},BBox_{2}^{tr},\dots,BBox_{n}^{tr}]\leftarrow$ BBoxes from $\hat{\omega}_{tr}^{t-1}$
+2:  $[BBox_{1}^{det},BBox_{2}^{det},\dots,BBox_{m}^{det}]\leftarrow$ BBoxes from $\hat{\omega}_{det}^{t-1}$
+3:  for  $BBox_{j}^{det}$ in $[BBox_{1}^{det},\dots,BBox_{m}^{det}]$ do
+4:     for  $BBox_{i}^{tr}$ in $[BBox_{1}^{tr},\dots,BBox_{n}^{tr}]$ do
+5:        Compute $IOU(BBox_{j}^{det},BBox_{i}^{tr})$
+6:        if $IOU>\delta_{iou}$ then
+7:           Discard $BBox_{j}^{det}$
+8:           break
+9:        end if
+10:     end for
+11:  end for
+12:  $\omega_{\det_{f}}^{t-1}\leftarrow$ Remaining detect queries assignment
 The initial design of the TBSP was intended merely to equip a detection network with basic tracking functionality.
@@
 Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training.
-This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the model to optimize itself.
-For a discussion on why TBSP is effective, please refer to the appendix.
+This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the method to optimize itself.
+Formally, we denote the predictions of track queries as $\widehat{Y}_{tr}$ and predictions of detect as $\widehat{Y}_{\text{det }}.Y_{\text{new }}$ is the ground-truths of newborn objects.
+The label assignment results for track queries and detect queries can be written as $\omega_{tr}$ and $\omega_{\text{det}}$.
+For frame $i$, label assignment for detect queries is obtained from bipartite matching among detect queries and newborn objects.
 We know that both track query and detect query contain information about the object’s bounding box (BBox).
 Let’s denote the BBox in $\hat{\omega}_{tr}^{t-1}$ as $[BBox_{1}^{tr}$, $BBox_{2}^{tr},...,BBox_{n}^{tr}]$, where $n$ is the number of track queries at time $t-1$.
-Each query in $\hat{\omega}_{tr}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
+Each query assignment in $\hat{\omega}_{det}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
 For each element in $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$, calculate the Intersection over Union (IOU) with each element in $[BBox_{1}^{tr},BBox_{2}^{tr},...,BBox_{n}^{tr}]$.
@@
 $\hat{\omega}_{tr}^{t}\,\,=\omega_{det_{f}}^{t-1}\cup\,\,\hat{\omega}_{tr}^{t-1}\,\,$ (1)
-The feasibility of TBSP is illustrated in the Fig.
-4.
+The feasibility of TBSP is illustrated in the Algorithm.
+1.
 Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted.
-Simultaneously, to better illustrate the performance of our model compared to MOTR, we have not adopted the labeling strategy used in CO-MOT[8] and MOTRv3[9].
-3.3 Loss and Training Methods
-Due to MOTR’s similarity to RNN in its inference approach, its batch size is limited to 1.
-This often results in suboptimal utilization of GPU resources during training.
-Additionally, MOTR learns both image and motion features simultaneously.
-To expedite model training, this paper adopts a three-stage training strategy.
+Simultaneously, to better illustrate the performance of our method compared to MOTR, we have not adopted the labeling strategy used in CO-MOT [7] and MOTRv3 [8].
+Figure 5: Refinement process for DecoderTracker on partially trained track queries.
+When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized.
+At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Figure.
+1.
+The process in DecoderTracker+ is similar, but there are some differences in the removal and addition of track queries.
+The detailed operations are discussed in the subsequent sections.
+3.1.4 Query and Query Positions
+MOTR was initially built upon Deformable DETR [11].
+The original DETR was considered slow to converge due to its Hungarian matching algorithm.
+Although Deformable DETR significantly improved the convergence speed of DETR-like methods through architectural enhancements, subsequent works such as DN-DETR [30] and DAB-DETR [31] further corroborated the assertion that the Hungarian matching algorithm was indeed problematic.
+In DETR and Deformable DETR, the queries are typically initialized as a set of non-interpretable parameters, which is evidently suboptimal for training.
+Drawing inspiration from RT-DETR’s query generation approach, we directly generate queries from the input decoder features f.
+This process can be specifically represented as follows:
+$q_{d}&=\text{LayerNorm}(\text{Linear}(f))\\ \end{aligned}$ (2)
+One more thing, to expedite method convergence, we implement a Query Positions generation strategy akin to CO-MOT.
+Drawing inspiration from DAB-DETR [31], we employ a fixed transformation to convert reference points into Query Positions.
+While experimentation with the RT-DETR approach involving a Multilayer Perceptron (MLP) to transform reference points shows no significant performance discrepancy, it is important to note that the use of MLP substantially increases training time.
+3.1.5 Loss and Training Methods
+To expedite method training, this paper adopts a three-stage training strategy.
 In the first stage, emphasis is placed on training the network to extract target image features efficiently.
-The second stage involves training the model on weak tracking aspects using TBSP as a pretraining step for the third stage.
-Lastly, the third stage employs the TALA strategy to comprehensively enhance the model’s tracking performance.
+The second stage involves training the method on weak tracking aspects using TBSP as a pretraining step for the third stage.
+Lastly, the third stage employs the TALA strategy to comprehensively enhance the method’s tracking performance.
 In the initial stage, the network is trained as a detection network, and the loss function is formulated defined as Eq.3:
-$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)}{V_{i}}$ (2)
-$\mathcal{L}$ is the loss of single frame.
-Notably, during this stage, we employed the IoU-aware Query Selection from RT-DETR, so the loss function utilized is from RT-DETR.
-Unlike the original DETR, this stage introduces a loss termed $\mathcal{L}_{enc}$.
-This loss involves treating the initialized queries as boxes for the purpose of loss computation.
-The specifics are outlined below:
-$\mathcal{L}_{enc}(\hat{y},y)&=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},\hat{b},y,b)\\ &=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},c,IoU)\end{aligned}$ (3)
-where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively.
-Finally, the overall loss is as follows:
 $\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (3)
-where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[27], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[28], $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients.
+where $\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the frame predictions.
+The loss is calculated by ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$.
+$\mathcal{L}_{cls}$ denotes the focal loss[32], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[33].
+$\mathcal{L}_{enc}$ is the anchor loss of DAB-DETR [31] and $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients, which are set to 1,2,5,5 in our method, just like DAB-DETR.
 In this stage, the network is capable of initially learning the appearance features of objects to be tracked.
 During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time.
-Additionally, various data augmentation techniques [29, 30, 31] can be employed to expedite the model’s extraction of appearance features.
+Additionally, various data augmentation techniques [34, 35, 36] can be employed to expedite the method’s extraction of appearance features.
 The second stage of training is conducted using MOTR’s Collective Average Loss.
@@
 The loss functions used in this stage and the third stage are as Eq.4.
-$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\sum_{n=1}^{N}{\left(\mathcal{L}\left(\left.\widehat{Y}_{tr}^{t}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)+\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)\right)}}{\sum_{n=1}^{N}{\left(V_{t}\right)}},N=5$ (5)
+[…]
 where $V_{i}=V_{tr}^{i}+V_{det}^{i}$ denotes the total number of ground-truths objects at frame $i$.
 $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively.
+$\mathcal{L}$ is the loss of single frame, which is similar to Eq.3.
 At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects.
 Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage.
+Thanks to our various strategies and improvements to the network, we achieved huge improvements in training time and inference speed compared to MOTR.
 The third stage of training involves the utilization of TALA during the training phase, while simultaneously deactivating the TBSP.
 Experimental evidence indicates that if the training epochs during this stage are sufficient, TBSP can also be disabled during the inference phase without significantly compromising performance.
-This finding underscores the robustness of the model, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
+This finding underscores the robustness of the method, suggesting that once adequately trained in the third stage, TBSP can be turned off during inference with negligible impact on overall performance.
+3.2 DecoderTracker+
+Following the improvements in previous sections that led to the proposal of DecoderTracker, we observed suboptimal enhancements in inference speed, failing to meet anticipated performance benchmarks.
+Through latency profiling of individual modules, we identified the decoder as the primary bottleneck limiting acceleration.
+Subsequent in-depth analyses of GPU memory utilization and query quantity revealed that the core issue likely stems from dynamic data structures.
+Compared to static data configurations, dynamic data introduces inference challenges such as inefficient memory allocation scheduling, with empirical evidence provided in Figure 6.
+To address this, we present DecoderTracker+ in this chapter.
+By implementing a Fixed-Size Query Memory(FSQM) mechanism that maintains a constant number of tracking queries, our approach enables deeper optimization compatibility with modern deep learning frameworks ,e.g., PyTorch, TensorFlow.
+Furthermore, the deterministic data structure facilitates straightforward conversion to TensorRT, paving the way for accelerated deployment of end-to-end MOT methods on edge devices.
+DecoderTracker+ maintains the core architectural framework and operational workflow of its predecessor, as systematically illustrated in Figure 2, while introducing strategic innovations primarily in two key dimensions: Fixed-Size Query Memory and Attention Computation Optimization.
+Subsequent subsections provide detailed technical expositions of these advancements.
+Figure 6: The two figures illustrate the changes in Decoder memory occupancy and the latency of various components for MOTR, its base model Deformable DETR, our DecoderTracker, and DecoderTracker+ on the MOT17-14 sequence.
+In the right figure, the backbone of our methods is represented by FENet, and since they do not have an Encoder, their latency on the Encoder is 0.
+The memory changes of DecoderTracker are basically consistent with those of MOTR, so they are not displayed here.
+It can be clearly seen from the figures that dynamic data leads to unstable memory occupancy of the Decoder and abnormally high latency.
+3.2.1 Fixed-Size Query Memory
+FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$, where $N$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension.
+Compared to dynamic query mechanisms, this approach introduces:
+$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N}}_{\text{confidencevector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (5)
+to model memory for the continuous tracking process.
+During initialization, set $q_{ftr}^{\mathcal{M}}=\mathbf{0}$ and $C_{\text{tr}}^{(0)}=\mathbf{0}$ for all queries, where all queries are initialized to zero to avoid interference with networks that do not rely on attention mechanisms.
+Additionally, all inactive queries are also initialized to zero.
+Moreover, a global ID pool, denoted as $\mathcal{ID}$, is introduced.
+Each active tracking query is assigned a unique ID that accompanies the query as it moves across video frames.
+When a query is no longer associated with a tracked object, its corresponding ID is deleted and will not be reused.
+A new ID is assigned to new active queries.
+By the way, all inactive queries are assigned an ID of -1 and their bounding boxes are initialized with zero values.
+During each frame processing, iterative updates are performed as follows:
+  1. 1.
+    New query injection: For detection queries assignment $\{\hat{\omega}_{\text{det}}^{(i)}\}$ with detection confidence $c_{i}>\tau_{\text{in}}$, each query is directly assigned the first available ID from the global ID pool (which initially contains the value -1 for new queries).
+The corresponding query in the memory pool is updated using a momentum update rule:
+    $j=\mathop{\arg\min}\limits_{k}(\mathcal{ID}[k]=-1),$ (6)
+    $\hat{\omega}_{tr}^{\mathcal{M}}{(t)}[j]=\hat{\omega}_{\text{det}}^{(i)}$
+    After updating the query, its corresponding ID is assigned from the global pool.
+  2. 2.
+    Query removal: For the existing queries in the memory pool, confidence is monitored over time.
+If a query’s confidence falls below a threshold $\tau_{\text{out}}$ for three consecutive frames, it is considered for removal.
+Upon removal, the query is reset to zero, and its corresponding ID is set to -1.
+Specifically:
+    $\text{If }C_{\text{tr}}^{(t-3)}[n]<\tau_{\text{out}},C_{\text{tr}}^{(t-2)}[n]<\tau_{\text{out}},C_{\text{tr}}^{(t-1)}[n]<\tau_{\text{out}},$ (7)
+    then the query is removed by setting:
+    $q_{tr}^{\mathcal{M}}[n]=\mathbf{0},\quad\mathcal{ID}[n]=-1.$ (8)
+    where $S(\cdot)$ represents the prediction score function. The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
+It is also important to emphasize that, aside from this, the processing of queries in FSQM is consistent with that in DecoderTracker.
+The specific process will not be repeated here, but can be referred to in the previous sections.
+Algorithm 2 FSQM Online Update Procedure
+0:  New query set $\hat{\omega}_{\text{det}}=\{q_{i},c_{i},BBox_{i}\}_{i=1}^{M}$, memory pool $\hat{\omega}_{tr}^{\mathcal{M}}=(q_{\text{tr}},C_{\text{tr}},\mathcal{ID},BBox_{tr})$
+0:  Updated memory pool $\mathcal{M}^{\prime}$
+1:  $\hat{\omega}_{\text{valid}}\leftarrow\{\hat{\omega}_{i}|c_{i}>\tau_{\text{in}}\}$Filter valid new queries}
+2:  if $\hat{\omega}_{\text{valid}}\neq\emptyset$ then
+3:     for $q_{j}\in\mathcal{\hat{\omega}}_{\text{valid}}$ do
+4:        Assign the first available ID from the global ID pool $\mathcal{ID}$ to $q_{j}$
+5:        $k\leftarrow\mathop{\arg\min}\limits_{n}\mathcal{ID}[n]=-1$Find the first inactive query}
+6:        $q_{\text{tr}}[k]\leftarrow q_{j},BBox_{\text{tr}}[k]\leftarrow BBox_{j}$Momentum update}
+7:        $C_{\text{tr}}[k]\leftarrow c_{i}$Reset confidence to 1.0}
+8:     end for
+9:  end if
+10:  for $n\in\{1,...,N\}$ do
+11:     if $C_{\text{tr}}[n]<\tau_{\text{out}}$ for 3 consecutive frames then
+12:        $q_{\text{tr}}[n]\leftarrow\mathbf{0},BBox_{\text{tr}}[k]\leftarrow\mathbf{0}$Set query and BBox to zero (inactive)}
+13:        $\mathcal{ID}[n]\leftarrow-1$Reset ID to -1}
+14:     end if
+15:  end for
+16:  return $\mathcal{M}^{\prime}=(Q_{\text{tr}},C_{\text{tr}},\mathcal{ID})$
+3.2.2 Attention Mechanism Optimization
+Since some queries in FSQM are empty, we need to explore the network structure to ensure that these empty queries do not affect the final results.
+The network processing these queries consists only of linear layers, normalization layers, dropout, activation function, cross-attention with image features, and finally, self-attention.
+Our first step is to analyze how empty queries may impact the output of these networks.
+Firstly, we will analyze some unaffected networks.
+These network architectures inherently avoid commingling active and inactive queries in their computational processes.
+Therefore, during inference, there is no risk of interference arising from their interactions.
+Our subsequent analysis need only focus on verifying whether these networks produce gradients during the training.
+For linear layers, zero-valued queries produce the following output:
+$W\mathbf{0}+b=b$ (9)
+Here, $W$ is the weight matrix, $\mathbf{0}$ is the zero-valued query, and $b$ is the bias term.
+Since the zero-valued query contributes nothing, only the bias remains, which can be neutralized through layer normalization.
+After applying layer normalization, the output is given by:
+$\hat{y}=\frac{y-\mu}{\sigma}\cdot\gamma+\beta$ (10)
+where $y=b$, $\mu$ and $\sigma$ are the mean and standard deviation of the output across the batch, and $\gamma$ and $\beta$ are learnable parameters.
+Since the bias term $b$ is small in practice, empirical tests show that its $L_{2}$ norm $\|b\|_{2}$ is very small, with $\|b\|_{2}<0.03$ in our implementation.
+For normalization layers, since the zero-valued queries do not produce any meaningful activation, they do not affect the overall normalization process.
+The query’s zero value is normalized along with other queries, resulting in minimal influence on the normalized output.
+This can be expressed as:
+$q_{i}=\frac{q_{i}-\mu}{\sigma}\cdot\gamma+\beta\quad\text{for active queries }q_{i}$ (11)
+where $\hat{q}_{i}$ is the normalized output for an active query $q_{i}$. For inactive queries (those initialized to zero), the effect is negligible due to their zero contribution.
+For dropout, which randomly sets a fraction of the inputs to zero during training, inactive queries do not contribute to the dropout process.
+Inactive queries, having values of zero, remain unaffected by dropout, and the overall output for those queries is still zero.
+Mathematically, dropout can be expressed as:
+$\text{Dropout}(\omega_{i})=\begin{cases}0&\text{if query is inactive (}q_{i}=0\text{)}\\ q_{i}\cdot d&\text{if query is active}\end{cases}$ (12)
+where $d\in[0,1]$ is a random variable representing the dropout rate. Inactive queries (zero-valued) remain zero regardless of whether dropout is applied.
+For activation function, since only GELU is used to process queries, we will only discuss GELU here. In the case of GELU, the output for a zero-valued query is:
+$\text{GELU}(q_{i})=0.5\omega_{i}\left(1+\tanh\left(\sqrt{\frac{2}{\pi}}(q_{i}+0.044715q_{i}^{3})\right)\right)\quad\text{for }q_{i}=0\implies\text{GELU}(0)=0.$ (13)
+Thus, for activation function, zero-valued queries result in zero output, ensuring that they do not influence the subsequent layers.
+For cross-attention, given key matrix $K$ and dimension $d$ of Query, the attention scores become:
+$\text{Attn}(Q_{\text{zero}},K)=\frac{\mathbf{0}\cdot K^{T}}{\sqrt{d}}=\mathbf{0}$ (14)
+resulting in zero contribution after softmax normalization.
+In summary, for linear, normalization, dropout, activation layers and cross-attention, zero-valued queries contribute nothing to the computation, ensuring that they do not interfere with the active queries during the forward and zero-valued in the forward inherently prevent gradient computation, thereby eliminating any impact on the training.
+Regarding the self-attention layers, since the queries, keys, and values used in computing self-attention are all derived from the queries themselves, inactive queries inevitably introduce interference to the output.
+While a detailed analysis of this phenomenon is omitted here, we address this issue effectively by implementing a masking mechanism.
+Here we use the self attention layer in Decoder to illustrate, and the similarity in TAN will not be further elaborated.
+When calculating self-attention, we first compute an attention weight matrix $A$:
+$A=\text{softmax}\left(\frac{q(q^{\prime})^{T}}{\sqrt{d}}+\log(\text{Mask})\right)$ (15)
+where $q\in\mathbb{R}^{2N\times d}$ is obtained by concatenating $q_{tr}^{\mathcal{M}}$ from FSQM and $q_{det}$, i.e., $q=\text{concat}(q_{tr}^{\mathcal{M}},q{det})$ in first decoder layer, while in other layers, $q$ is the output from the previous layer.
+$(\text{Mask}\in\mathbb{R}^{2N\times 2N}$ is the mask we need to generate.
+Its dimensions are the same as those of $A$, and it ensures that after applying the softmax function, the unwanted values are set to zero.
+By appropriately using this mask during the final self-attention output $A\times V$ where $V$ is derived from $Q$ with positional encoding, the effect of empty queries on the final result can be effectively eliminated.
+Here, we extend $\mathcal{ID}$ and then generate the mask.
+$\mathcal{ID}_{\text{ext}}=\text{Concat}(\mathcal{ID}_{\text{tr}},\mathbf{0}_{N})$ (16)
+where, $\mathbf{0}_{N}$represents a fixed identifier for detect queries.
+$\text{Mask}_{ij}=\begin{cases}0&\text{if }\mathcal{ID}_{ext}[i]=-1\lor\mathcal{ID}_{ext}[j]=-1\\ 1&\text{otherwise}\end{cases}$ (17)
+By implementing this mask mechanism, we ensure that values associated with inactive queries in matrix $A$ are nullified to zero, thereby eliminating their influence on the final output.
+Notably, as demonstrated through our preceding analyses, the inactive query positions inherently maintain zero values across all decoder layers.
+Consequently, this masking mechanism uniformly applies to self-attention operations in every decoder layer without requiring layer-specific adaptations.
+In terms of the mask in the self-attention layer of the TAN, the generation process and mechanism are largely consistent with those described earlier; however, the dimensions in TAN are fewer, and there is no need to expand $\mathcal{ID}$.
 4 Experiments
 4.1 Datasets and Metrics
-To validate DecoderTracker’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[32], MOT17[33], and KITTI[34].
+To validate DecoderTracker and DecoderTracker+’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[37], MOT17[38], and KITTI[39].
 DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures.
@@
 The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[40, 41].
-These metrics collectively provided a comprehensive and robust evaluation of our model’s tracking performance.
+These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance.
 HOTA assessed the tracking accuracy based on the spatial and temporal overlap between predicted and ground truth tracks.
@@
 MOTA offered an overall evaluation of tracking accuracy, considering false positives, false negatives, and identity switches.
-These metrics formed the basis for a thorough evaluation of our model’s tracking capabilities.
+These metrics formed the basis for a thorough evaluation of our method’s tracking capabilities.
 4.2 Implementation Details
 The rationale for adopting DecoderTracker originates from our observation of the high training costs and suboptimal inference speed associated with MOTR.
-Thus, our aim is to propose a new baseline model surpassing MOTR’s performance.
+Thus, our aim is to propose a new baseline method surpassing MOTR’s performance.
 To ensure a fair comparison, we exclusively employ methods enhancing tracking performance, while omitting improvement techniques utilized in MeMOTR [9], MOTRv2 [18], MOTRv3 [8], and CO-MOT [7].
@@
 Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU.
-One more thing, he image size for the input network is $640\times 640$.
-During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti.
-Subsequent stages utilized a uniform batch size of 1.
-For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [37], employing random displacement to generate video segments with pseudo-trajectories.
-Training comprised 120 epochs in the first stage, followed by 30 epochs in the second.
-The third stage extended to 20 epochs with TBSP, or 55 without.
-For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP.
-For KITTI, initial training involved 80 epochs, followed by 25 in the second stage.
-The third stage consisted of 30 epochs with TBSP or 70 without.
-Additionally, approximately 5k images from BDD100k [38] were selected for additional training.
-4.3 Performance Evaluation and Comparison
-In Table 1, a comprehensive comparison of training time and speed was conducted between DecoderTracker and the MOTR series.
-It is crucial to note that the datasets used for comparing training time and inference speed exhibit variations.
-This disparity primarily arises from the act that the publicly disclosed training times and inference speeds of the MOTR series are based on different datasets.
+One more thing, the image size for the input network is $640\times 640$.
 (a) Comparison of Training Time on MOT17
@@
 DecoderTracker 19.6
+DecoderTracker+ 28.8
 Table 1: Comparison of training time (including all previous stages) and speed between DecoderTracker and MOTR series on a single V100 GPU.
-The data for the MOTR series is sourced from [3], [9] or their open source address.
-DecoderTracker demonstrates a substantial advantage over the MOTR series, particularly in terms of training time and speed.
-The primary reason behind this efficiency gain lies in the architectural simplicity of DecoderTracker.
-Unlike MOTR, DecoderTracker adopts a streamlined structure that eliminates the need for an encoder, resulting in a more computationally frugal design.
-This reduction in complexity enables DecoderTracker to achieve training completion within 68.7 hours, utilizing only 1 Nvidia GeForce 2080ti GPU.
-In contrast, MOTR relies on 8 Nvidia GeForce 2080ti GPUs and takes about 96 hours for training.
-The absence of an encoder in DecoderTracker contributes to its swift training process, emphasizing the model’s efficiency in resource utilization.
-Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[32] benchmark.
+The training time of DecoderTracker+ is basically the same as that of DecoderTracker, so it will not be listed separately here.
+Figure 7: IDF1-MOTA-HOTA and IDF1-MOTA-FPS comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on Dancetrack test sets.
+For the left image, the horizontal axis is MOTA, the vertical axis is IDF1, and the radius of the circle is HOTA.
+For the right image, the horizontal axis is FPS, the vertical axis is IDF1, and the radius of the circle is HOTA.
+”*” represents additional data was used for training.
+Our DecoderTracker+* achieves the best IDF1 and HOTA and comparable speed performance.
+Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[37] test set.
 Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR.
-Methods HOTA IDF1 DetA AssA MOTA
+* means that extra data was used during training.
+Methods IDF1 HOTA AssA DetA MOTA
 Non-End-to-end:
-ByteTrack[14] 47.3 52.5 71.6 31.3 89.5
-GTR[39] 48.0 50.3 72.5 31.9 84.7
-DST-track[40] 51.9 51.0 82.3 34.6 84.9
-OC-SORT[41] 54.1 54.2 80.4 38.0 89.4
+TransTrack[15] 45.2 45.5 27.5 75.9 87.7
+GTR[21] 50.3 48.0 31.9 72.5 84.7
+DST-track[42] 51.0 51.9 34.6 82.3 84.9
+ByteTrack[14] 52.5 47.3 31.3 71.6 89.5
+OC-SORT[22] 54.2 54.6 38.0 80.4 89.4
+Deep OC-SORT[5] 61.5 61.3 45.8 82.2 92.3
 End-to-end:
-MeMOTR[10] 63.4 65.5 77.0 52.3 85.4
-MOTR[3] 54.2 51.5 73.5 40.2 87.6
-DecoderTracker(Ours) 54.2 54.2 78.7 37.4 87.8
-On the Dancetrack test benchmark, DecoderTracker(TBSP) achieves a speed of 18.5 frames per second (FPS) using a single V100, while the MOTR series attains speeds of 9.5 FPS (MOTR), 6.9 FPS (MOTRv2), and 10.6 FPS (MOTRv3).
-Particularly noteworthy is the remarkable 19.6 FPS achieved by DecoderTracker, emphasizing its outstanding performance in inference speed.
-Table 2, Table 3, and Table 4 exhibit the final results of end-to-end DecoderTracker on the Dancetrack, MOT17, and KITTI benchmark, respectively.
-In contrast to MOTR, DecoderTracker demonstrates competitive performance across multiple metrics on both datasets.
-It is important to highlight that the performance of MeMOTR is provided in Tables 2 and 3.
-The inclusion of MeMOTR’s performance aims to underscore the considerable potential for advancement and enhancement in both MOTR and our DecoderTracker.
-MeMOTR[10] integrates additional modules and performance-enhancing techniques, along with supplementary datasets during training.
-Consequently, comparing its tracking performance directly with that of DecoderTracker and MOTR may not be entirely fair.
-Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17[33] set.
-method HOTA IDF1 DetA AssA MOTA
+MOTR[2] 51.5 54.2 40.2 73.5 79.7
+DecoderTracker 54.2 54.2 35.7 78.7 87.8
+DecoderTracker+ 54.4 54.5 35.6 79.0 88.1
+DecoderTracker* 62.9 60.8 47.1 81.2 91.1
+DecoderTracker+* 62.2 61.2 46.0 81.8 91.9
+Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [38] set. * means that extra data was used during training.
+method IDF1 HOTA DetA AssA MOTA
 Non-End-to-end:
-FairMOT[42] 59.3 72.3 60.9 58.0 73.7
-ByteTrack[14] 62.8 77.2 63.8 62.2 78.9
-StrongSORT[39] 63.5 78.3 63.6 63.7 78.5
-UTM[43] 64.0 78.7 65.9 62.5 81.8
-OC-SORT[41] 63.2 77.5 63.2 63.2 78.0
-Deep OC-SORT[6] 64.9 80.6 64.1 65.9 79.4
+YOLOTracker[43] 65.1 - - - 71.9
+FairMOT[44] 72.3 59.3 60.9 58.0 73.7
+MSPNet[45] 74.6 - - - 71.9
+ByteTrack[14] 77.2 62.8 63.8 62.2 78.9
+OC-SORT[22] 77.5 63.2 63.2 63.2 78.0
+StrongSORT[21] 78.3 63.5 63.6 63.7 78.5
+UTM[46] 78.7 64.0 65.9 62.5 81.8
+Deep OC-SORT[5] 80.6 64.9 64.1 65.9 79.4
 End-to-end:
-MeMOTR[10] 58.8 69.0 59.6 58.4 72.5
-MOTR[3] 57.2 59.4 58.9 55.8 71.9
-DecoderTracker(Ours) 55.7 70.5 55.7 56.0 66.5
-For Dancetrack (Table 2), DecoderTracker achieves comparable results to MOTR, with same scores in HOTA but outperforming MOTR in DetA.
-Specifically, DecoderTracker achieves an impressive MOTA, indicating its overall superior performance in multi-object tracking tasks.
-In the MOT17 (Table 3), DecoderTracker demonstrates competitive performance, particularly in handling complex tracking scenarios, with competitive scores in HOTA and AssA.
-Despite trailing slightly behind MOTR in these metrics, DecoderTracker compensates with a comparable DetA and excels in IDF1 by almost 10 points.
-This advantage can be attributed to DecoderTracker’s simplified network structure, which enables competitive tracking accuracy with reduced training data demand.
-Additionally, DecoderTracker maintains robust overall performance, as evidenced by its high MOTA score, highlighting its suitability for real-world tracking applications.
+MOTR[2] 59.4 57.8 60.3 55.7 73.4
+DecoderTracker 70.5 55.7 55.7 56.0 66.5
+DecoderTracker+ 70.4 55.8 55.2 55.8 66.7
+DecoderTracker* 77.2 61.8 62.1 60.1 76.8
+DecoderTracker+* 77.5 62.2 62.5 60.7 76.7
+During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1.
+For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [47], employing random displacement to generate video segments with pseudo-trajectories.
+Training comprised 120 epochs in the first stage, followed by 30 epochs in the second.
+The third stage extended to 20 epochs with TBSP, or 55 without.
+And, to address the issue of the limited scale of the MOT17 dataset, this paper employs a hybrid dataset.
+It integrates the MOT20 dataset and the complete CrowdHuman dataset, which has been spatially shifted, to expand data diversity.
+The amount of new data significantly exceeds the benchmark set (by approximately 4.3 times).
+To balance training bias, a stratified sampling mechanism was designed: within each epoch, samples from the MOT17 dataset are resampled twice, while samples from the supplementary sets are sampled once.
+This approach not only alleviates data skewness but also enhances the learning of benchmark set features, ensuring that the model maintains its original performance advantages while improving its adaptability to dense scenes.
+For Dancetrack, ensuring parity without extra data like MOTRv2, the first stage was 15 epochs, the second was 9, and the third, if TBSP was used, was 9 epochs, or 18 without TBSP.
+To better compare with the state-of-the-art (SOTA) methods, when testing on the test set, we also used the CrowdHuman dataset as supplementary data and the validation set of Dancetrack was also used.
+For KITTI, initial training involved 80 epochs, followed by 25 in the second stage.
+The third stage consisted of 30 epochs with TBSP or 70 without.
+Additionally, approximately 5k images from BDD100k [48] were selected for additional training.
+4.3 Main Result
+The comprehensive benchmarking in Table 1 demonstrates DecoderTracker’s dual superiority in training efficiency and inference speed over MOTR-series methods.
+This design achieves 68.7 training hours on a single GPU, equivalent to 7.1× higher GPU-hour efficiency versus MOTR’s 768 GPU-hour consumption (8 GPUs × 96 hours).
+The phased training regimen—12.1 hours for feature embedding, 30.8 hours for temporal modeling, and 43.4 hours for TBSP-enhanced prediction—systematically constructs tracking capabilities while maintaining 83% shorter wall-clock time than competitors.
+Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS,
+For DanceTrack ,the experimental results shown in Table 2 reveal that DecoderTracker and its enhanced variant DecoderTracker+ achieve comparable precision levels across core evaluation metrics.
+Both models demonstrate superior detection accuracy over existing end-to-end methods, with DetA scores of 78.7 and 79.0 respectively, significantly outperforming MOTR’s 73.5.
+Their MOTA metrics (87.8 and 88.1) approach SOTA non-end-to-end methods like ByteTrack (89.5), while maintaining balanced HOTA performance at 54.2 and 54.5.
+This parity suggests that architectural refinements between the two variants primarily yield marginal improvements rather than fundamental performance shifts.
+Notably, both configurations exhibit identical IDF1 scores (54.2 and 54.4), confirming consistent identity preservation capabilities.
+The minimal performance delta between variants implies optimized parameter utilization rather than architectural overhauls.
+When extended with supplementary training data (DecoderTracker* variants), both models achieve substantial gains, attaining 91.1-91.9 MOTA and 60.8-61.2 HOTA.
+While their AssA scores (35.6-35.7) remain below MOTR’s 40.2, the detection-centric advantages manifest through DetA superiority over all end-to-end and most non-end-to-end competitors.
+These findings position the DecoderTracker framework as a robust baseline for end-to-end tracking, achieving detection reliability comparable to sophisticated multi-stage systems while preserving model efficiency.
+In the MOT17 benchmark (Table 3), DecoderTracker shows strong performance, especially in complex tracking scenarios.
+While slightly behind MOTR in HOTA and AssA, it compensates with comparable DetA and excels in IDF1 by nearly 10 points.
+This advantage is due to its simplified network structure, which achieves competitive accuracy with less training data.
+Moreover, DecoderTracker maintains robust overall performance, as shown by its high MOTA score, making it suitable for real-world applications.
+When additional training data is introduced, DecoderTracker’s performance further improves.
+For instance, DecoderTracker* achieves a significant boost in IDF1 to 77.2, HOTA to 61.8, and MOTA to 76.8.
+Similarly, DecoderTracker+* also shows substantial gains.
+This clearly demonstrates the positive impact of more training data on DecoderTracker’s tracking capabilities, highlighting its potential for even better performance with increased data support.
 Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI [39] set.
@@
 DecoderTracker 51.5 56.8 58.4 72.2 83.2 73.8
+DecoderTracker+ 51.1 56.4 58.8 72.6 82.1 73.9
 Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark.
@@
 These results position DecoderTracker as a strong contender in MOT on the KITTI.
-4.4 Ablation Study
+Figure 8: The visualisation results of our proposed DecoderTracker on various datasets show that DecoderTracker performs well both under long time series and short time series, and also has good tracking performance for different kinds of objects.
+The visualization results of DecoderTracker and DecoderTracker+ are essentially identical, so we won’t display the results of D+ here.
+4.4 Ablation Study Of DecoderTracker
 To assess the proposed training strategy and the efficacy of TBSP, ablation experiments were conducted on the Dancetrack benchmark.
 The training set includes the entire Dancetrack training data, with testing involving 25 sequences from the validation set.
-Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR[48] or employing an MLP, on the model’s performance was investigated.
+Additionally, the impact of two methods for generating query positions—utilizing the sinusoidal functions for encodings from DAB-DETR [31] or employing an MLP, on the method’s performance was investigated.
 (a) Performance of DecoderTracker at different training stages
-Training Stage HOTA AssA MOTA
-First 38.5 21.6 76.8
-Second 44.5 25.5 84.2
-Final (TBSP) 51.9 34.6 84.9
-Final 53.6 35.9 84.7
-(b) The presence or absence of TAN during the second training stages can significantly impact the performance of the model.
-Model HOTA AssA MOTA
-Second 44.5 25.5 84.2
-Second(w/o) 42.1 21.4 80.8
-Final 53.6 35.9 84.7
-Final(w/o) 53.4 36.1 84.5
-(c) The impact of $\delta_{iou}$ on model performance when using TBSP
-$\delta_{iou}$ HOTA AssA MOTA
-0.3 49.7 32.3 79.1
-0.5 51.9 34.6 84.9
-0.7 51.3 34.8 83.2
-0.9 50.0 32.8 81.3
+Training Stage IDF1 HOTA MOTA
+First 38.6 38.5 76.8
+Second 44.3 44.5 84.2
+Final (TBSP) 51.7 51.9 84.9
+Final 53.4 53.6 84.7
+(b) The impact of $\delta_{iou}$ on method performance when using TBSP
+$\delta_{iou}$ IDF1 HOTA MOTA
+0.3 50.1 49.7 79.1
+0.5 51.7 51.9 84.9
+0.7 51.3 51.3 83.2
+0.9 49.8 50.0 81.3
+(c) The presence or absence of TAN during the second training stages can significantly impact the performance of the method.
+method IDF1 HOTA MOTA
+Second 44.4 44.5 84.2
+Second(w/o) 41.9 42.1 80.8
+Final 53.4 53.6 84.7
+Final(w/o) 53.2 53.4 84.5
 (d) Effect of query position generation strategy
-Setting HOTA MOTA FPS
-sin/cos 53.3 85.1 18.5
-MLP 53.6 84.7 19.6
+Setting IDF1 HOTA FPS
+sin/cos 85.1 53.3 18.5
+MLP 84.7 53.6 19.6
 Table 5: Ablation Studies on Our Proposed DecoderTracker on the DanceTrack validation set.
@@
 Position generation strategies, MLP and sinusoidal functions, yield similar performance in terms of detection accuracy, with slight differences in FPS.
-Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and model configurations, showcasing its adaptability and versatility in MOT field.
-It is noteworthy that in Table 5(b), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages.
+Overall, the ablation studies demonstrate the effectiveness of DecoderTracker across different training conditions and method configurations, showcasing its adaptability and versatility in MOT field.
+It is noteworthy that in Table 5(c), we demonstrate the impact of the presence or absence of TAN during the second phase of training on the network’s tracking performance at different stages.
 It can be observed that the presence or absence of TAN in the second phase has almost no effect on the final network performance.
 However, the absence of TAN does impact the tracking performance during the second phase, which also validates the effectiveness of our proposed TBSP strategy in enhancing tracking capabilities.
-5 Conclusion and Limitation
-This paper presents an efficient decoder-only end-to-end MOT (Multi-Object Tracking) model that boasts faster inference speeds and reduced training times.
-Additionally, this paper introduces new training strategies aimed at accelerating model convergence.
-Although device limitations prevented us from validating our training strategies on the original MOTR model, we believe these strategies are effective.
-This research provides a more efficient benchmark model for the multi-object tracking field compared to MOTR, laying a foundation for the emergence of real-time end-to-end MOT models and offering insights for future development.
-In subsequent work, we should further explore more efficient network architectures to construct real-time MOT models.
-Of course, we also plan to incorporate techniques from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to enhance tracking performance.
-Appendix A More Discusstion About TBSP
-Figure 5: Refinement process for DecoderTracker on partially trained track queries.
-When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized.
-At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Fig.
-1.
-Table 6: Comparison of DecoderTracker with and without using TBSP on the Dancetrack val set.
-Setting HOTA MOTA Epochs Total Time
-TBSP 53.6 84.7 27 56.7 hours
-w/o 53.4 84.2 40 85.2 hours
-The initial reason for developing the MO-YOLO model was not due to slow inference speed, but rather the extensive training resources required.
+4.4.1 More Discusstion About TBSP
+The initial reason for developing the DecoderTrack method was not due to slow inference speed, but rather the extensive training resources required.
 Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs.
 Each epoch took around 0.42 hours, with each GPU processing about 3,000 images per hour.
-Fortunately, our DecoderTracker model has addressed this issue.
+Fortunately, our DecoderTracker method has addressed this issue.
 Although its inference speed is not yet real-time, its convergence rate is already satisfactory.
@@
 Many examples demonstrate that undergoing weakly supervised training before strongly supervised training can enhance performance and speed up convergence.
-For instance, in the NLP field, large language models like BERT[49] and GPT[1] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns.
-Subsequently, these models undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering.
+For instance, in the NLP field, large language methods like BERT[53] and GPT[54] initially train on a vast amount of unlabeled or weakly labeled data to learn basic language patterns.
+Subsequently, these methods undergo strong supervision fine-tuning for specific tasks like sentiment analysis or question answering.
 This approach has been empirically proven to significantly boost the performance of LLMs.
+Table 6: Comparison of DecoderTracker with and without using TBSP on the Dancetrack val set.
+Setting IDF1 HOTA Epochs Total Training Time
+TBSP 54.2 53.6 27 56.7 hours
+w/o 53.5 53.4 40 85.2 hours
 In our field of computer vision, this technique is also quite common, with the most classic example being the SimCLR series[12, 13].
-Typically, these models are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
-We believe that the TBSP, a weak supervision process, enables the model to preliminarily learn object motion information that would otherwise require many more training iterations.
+Typically, these methods are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
+We believe that the TBSP, a weak supervision process, enables the method to preliminarily learn object motion information that would otherwise require many more training iterations.
 It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill.
 A similar situation may exist here.
-Although TBSP did not enhance model performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
-Appendix B Design of IMFENet
-The design of this feature network is intended to enhance the model’s efficiency and increase its inference speed.
-Thus, our design primarily references the network structure of YOLOv8.
-The backbone structure of our model is almost identical to that of YOLOv8.
-Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
+Although TBSP did not enhance method performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
+4.5 Ablation Study Of DecoderTracker+
+As can be seen from the results in Tables 2, 3, and 4, the performance of DecoderTracker and DecoderTracker+ is almost the same.
+This is not difficult to understand.
+In Section 3.3, we discussed that empty queries do not affect the network’s output results after reasonable mask generation.
+DecoderTracker+ does not improve the network; instead, it addresses the optimization issues caused by dynamic data.
+At the same time, to better illustrate the effectiveness of the generated masks, we provided the performance of DecoderTracker+ with and without masks on the Dancetrack val set in Table 7.
+The results show that when masks are not used, the inactive queries still have an impact on the final tracking results.
+Combining this with previous data, both theoretical and practical evidence demonstrates that the use of masks does not affect tracking performance when FSQM is increased after reasonable mask generation.
+Table 7: The impact of mask on the performance of DecoderTracker+ on the Dancetrack val set.
+mask IDF1 HOTA MOTA AssA
+53.6 53.4 84.9 36.0
+w/o 46.2 45.8 78.3 31.7
+4.5.1 Experimental Analysis of FENet
 Furthermore, we have made some modifications to the C2F module in YOLOv8, renaming our modified module to C2M.
@@
 These modifications are illustrated in Figure 9, and their impact on performance is shown in Table 8.
-Figure 6: The difference between our proposed model and YOLOv8 in terms of modules. Our modules are more lightweight. Specific model details can refer to the figure.3.
+Figure 9: The difference between our proposed method and YOLOv8 in terms of modules.
+Our modules are more lightweight.
+Specific method details can refer to the Figure.3.
+Additionally, it should be noted that there is not much difference between this part of YOLO11 and YOLOv8, so no comparison has been made
+Table 8: The impact of our network and YOLOv8 structure on DecoderTracker and DecoderTracker+ performance on the Dancetrack val set.
+Setting IDF1 HOTA MOTA AssA FPS
+DecoderTracker:
+FENet 53.8 53.6 84.7 35.9 19.6
+YOLOv8 51.4 52.4 84.2 36.1 18.9
+1.88 DecoderTracker+:
+FENet 53.6 53.4 84.9 35.7 28.2
+YOLOv8 52.1 53.2 85.4 36.8 26.7
+5 Conclusion and Limitation
 Figure 10: DecoderTracker and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 benchmark.
-Table 7: The impact of our network and YOLOv8 structure on DecoderTracker performance on the Dancetrack val set. Training time does not include the first stage
-Setting HOTA IDF1 MOTA AssA FPS
-IMFENet 53.6 53.8 84.7 35.9 19.6
-YOLOv8 52.4 51.4 84.2 36.1 18.9
-Appendix C Limitation
-Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection models.
-Both models fail under similar circumstances, as illustrated in Figure.7.
-This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the model.
+This paper proposes an efficient decoder-only end-to-end MOT method, which achieves faster inference speed and shorter training time compared to the baseline MOTR.
+Additionally, we introduce new training strategies aimed at accelerating the convergence of the method.
+Due to hardware limitations, we were unable to verify our training strategies on MOTR, but we believe these strategies are generalizable.
+Finally, we conducted a detailed and in-depth analysis of the slow inference process in MOTR, finding that dynamic data significantly impacts inference speed.
+Subsequently, we applied FSQM to address this issue, greatly enhancing the inference speed of the tracking-by-query end-to-end paradigm.
+Overall, this study provides a more efficient benchmark method for the multi-object tracking field, lays the foundation for the emergence of real-time end-to-end MOT methods, and offers insights for future development.
+In future work, we should further explore more efficient network architectures to build real-time MOT methods.
+Of course, we also plan to integrate the technologies from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to improve tracking performance.
+Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection methods.
+Both methods fail under similar circumstances, as illustrated in Figure.10.
+This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the method.
 Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR.
+Although we attempted to use a mixed dataset to address this issue, our method still underperforms compared to SOTA tracking-by-detection methods.
+How to resolve this issue will also be a key focus of future research.

### v6

@@
 However,the substantial computational resource consumption of these methods, coupled with the optimization challenges posed by dynamic data, results in less favorable inference speeds and training times.
-To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTrack, a novel end-to-end tracking method.
-Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTrack+ by incorporating a Fixed-Size Query Memory and refining certain attention layers.
-Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, with inference speeds 2.06 and 3.03 times faster than MOTR, respectively.
+To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTracker, a novel end-to-end tracking method.
+Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTracker+ by incorporating a Fixed-Size Query Memory and refining certain attention layers.
+Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, featuring a 2 to 3 times faster inference than MOTR, respectively.
+The proposed method is implemented in open-source code, accessible at https://github.com/liaopan-lp/MO-YOLO.
 keywords:
@@
 Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images.
-The success of transformer [1] methods in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT methods like MOTR [2] have been introduced in the MOT domain.
-These methods offer advantages over the widely used tracking-by-detection methods [3, 4, 5, 6], such as eliminating the need for manual design or selection of features.
-Additionally, some enhanced methods [7, 8, 9] of MOTR have demonstrated the potential of this method class in terms of tracking accuracy.
+The success of transformer [1] methods in various domains in recent years is well documented, and transformer-based end-to-end MOT methods such as MOTR [2] and MOTIP [3] have recently been introduced in the MOT domain.
+These methods offer advantages over the widely used tracking-by-detecting (TBD) methods [4, 5, 6, 7], such as eliminating the need for manual design or feature selection.
+Among them, the MOTR series based on the Tracking-by-Query(TBQ) paradigm extends object queries to track queries and achieves tracking by propagating these track queries across frames.
+This category of methods exhibits exceptional extensibility, with their core ideas frequently adopted by models in multiple domains such as 3D detection[8], motion prediction[9], and video understanding[10].
+In contrast, MOTIP[3], another end-to-end MOT method currently achieving higher accuracy, is somewhat inferior in this aspect.
+Its tracking mechanism is more akin to the TBD approach, except that its detection and tracking networks are jointly optimized during training.
+Thus, in-depth research on TBQ methods will not only significantly advance the development of MOT technology but also hold important implications for driving MOT toward greater efficiency and universality.
+Figure 1: Illustration of diverse MOT pipelines.
+For MOTIP and TBD, a detector first yields detections, followed by ID assignment networks or matching associators for tracking.
+In contrast, MOTR-series tracking-by-query methods use track queries to model targets and propagate them frame-by-frame.
+Our approach also employs tracking-by-query but features a simpler network; the proposed FSQM overcomes dynamic data challenges inherent in other tracking-by-query methods.
+However, the inference speed of these TBQ methods is often suboptimal.
+Their RNN-like linear inference structure makes them less GPU-friendly during training, and this linear training paradigm also renders data augmentation techniques commonly used in detection tasks inapplicable.
+Upon in-depth analysis, it becomes clear that the core mechanism of TBQ is accomplished by the decoder, modeling object trajectories and temporal correlations through query propagation and attention interactions, while the encoder merely handles static image feature extraction and contributes little to the core logic of dynamic tracking.
 […]
-Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios.
-The number of detect queries is typically fixed, while the number of track queries fluctuates throughout the tracking process.
-This discrepancy results in dynamic data during inference, which introduces optimization challenges that limit the actual runtime speed of the method.
-To address this, we propose the Fixed-Size Query Memory (FSQM), which pads the tracking query sequence with empty queries to maintain a consistent number of queries during inference.
-By eliminating dynamic data, we have developed DecoderTracker+, which has significantly improved inference speed compared to MOTR and DecoderTracker.
 At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the method to engage in less stringent learning might also be effective.
@@
 It is important to note that weak supervision, or self-supervised learning, has achieved significant success in many computer vision tasks [14, 15].
-Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during the training.
-Thus, DecoderTracker has been constructed by us.
+Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during training.
+Thus, we propose DecoderTracker, a highly efficient end-to-end TBQ tracking network.
+Its fundamental departure from prior approaches is visually discernible in Figure 1.
 After completing the aforementioned architectural optimizations mentioned above, we observed that while the training duration was significantly optimized, the inference speed did not improve as expected.
@@
 To address this issue, we proposed the Fixed-Size Query Memory (FSQM) for handling fixed-query tracking.
-By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTrack+.
-Experimental results demonstrated that, although DecoderTrack+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
+By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTracker+.
+Experimental results demonstrated that, although DecoderTracker+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
 In summary, the contributions of this paper are as follows:
@@
     The second key innovation involves the introduction of a unique training strategy characterized by a three-stage process.
-At the same time, we incorporated TBSP (Temporal Balanced Sampling and Propagation), a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods.
-These strategies accelerate method convergence, reduce overall training time, and enhance the efficiency of the training process.
+At the same time, we incorporated TBSP, a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods.
+These strategies accelerate method convergence, reduce overall training time by enhancing the efficiency of the training process.
   • 3)
-    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTrack+, which addresses the additional latency caused by dynamic data.
+    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTracker+, which addresses the additional latency caused by dynamic data.
 This provides a fully viable technical pathway for the engineering deployment of MOTR-like end-to-end MOT models.
@@
 2.1 MOTR Series
-Unlike tracking-by-detection methods exemplified by SORT [3, 14, 4], and previous transformer-based MOT methods [15, 16], MOTR [2] is the first method to achieve end-to-end multi-object tracking.
+Unlike tracking-by-detection methods exemplified by SORT [4, 16], and previous transformer-based MOT methods [17, 18], MOTR [2] is the first method to achieve end-to-end multi-object tracking.
 It extends the object query from DETR [19] into a track query, thereby enabling tracking capabilities.
@@
 This integration involves using YOLOX generated proposal boxes as anchors to initialize object queries, a method that not only bolsters detection accuracy but also helps to alleviate conflicts between detection and association tasks, all the while maintaining the query propagation features inherent to MOTR.
-Similarly, MeMOT [9] innovates by incorporating track queries that method and predict target trajectories iteratively across video frames.
+Similarly, MeMOTR [22] innovates by incorporating track queries that predict target trajectories iteratively across video frames.
 This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
-MOTRv3 [8] and CO-MOT [20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework.
-MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance.
-In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces ”shadows” to enhance the diversity and quantity of queries, further refining the MOT capabilities.
+MOTRv3 [23] and CO-MOT [24] focus on optimizing query interactions and label assignment within the MOTR framework.
+MOTRv3 adopts a release-fetch supervision strategy for balanced label assignments, while CO-MOT employs cooperative label assignment with ”shadows” to enhance query diversity.
+DNMOT[25] addresses severe occlusions through denoising training (simulating occlusion with noise) and a Cascaded Mask Module that coordinates query interactions to prevent suppression of occluded trajectories.
+DQFormer[26], meanwhile, uses Decoupled Query Augmentation to separate detection and association query branches, reducing feature conflicts.
+A
 2.2 Tracking by Detection
-Tracking by Detection are among the most classical and widely used types of MOT. These methods typically employ a detector to first identify objects and then use various other methods for tracking.
-In past years, the main methods within the Tracking by Detection category were SORT [3] and its derivative methods.
+Tracking by Detection frameworks are among the most classical and widely used approaches for MOT.
+These methods typically employ a detector to first identify objects and then use various other methods for tracking.
+In past years, the main methods within the TBD category were SORT [4] and its derivative methods.
 The primary approach of these methods is to use a Kalman Filter to predict the positions of targets in the next frame and then perform data association with the detection results provided by the detector.
-Many subsequent improvements to these methods [5, 14, 4, 21] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association.
-There are also a series of methods like OC-SORT [22] that purely improve the Kalman Filter-based tracking method using kinematic methods.
-In recent years, methods such as SUSHI [23], Glan [24] and PuTR [25] have emerged, which employ networks for tracking post-detection.
-The former two utilize Graph Neural Networks (GNN), while the last is based on transformer.
-2.3 Real Time Detection Method
+Many subsequent improvements to these methods [6, 16, 7] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association.
+There are also a series of methods like OC-SORT [27] and MotionTrack[5] that purely improve the Kalman Filter-based tracking method using kinematic methods.
+In recent years, within the TBD paradigm, a new wave of methods has emerged, such as SportMamba[28], TWiX[29], MSPNet[30] and Glan[31].
+These approaches leverage advanced neural network architectures—including spatiotemporal modeling networks and query-based attention mechanisms—to refine the post-detection tracking process, moving beyond the traditional reliance on Kalman filtering or standalone re-identification modules.
+Unlike earlier methods that primarily relied on kinematic predictions or appearance features for association, these newer frameworks integrate deeper spatiotemporal reasoning into the tracking pipeline, enabling more robust handling of complex scenarios such as long-term occlusion and fast motion.
 Figure 2: The overall structure of DecoderTracker.
@@
 This approach effectively addresses optimization issues caused by dynamic data.
+2.3 Real Time Detection Method
 When discussing real-time detection methods, most people initially think of YOLO [32], which transformed the problem of object detection into a regression method, achieving efficient object detection.
 The method divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions.
-The most advanced versions in the YOLO series currently are YOLOv8.2 [27] and YOLOv9 [28].
-Recently, the real-time detection method based on DETR, known as RT-DETR [29], has also demonstrated its extraordinary performance.
-RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the method structure.
-Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary.
-Therefore, inspired by the YOLO series [28, 27], we designed a feature extraction network.
+Among recent variants in the YOLO family, YOLOv8 and its upgraded version YOLO11[33] stand out for their superior stability and migration capability, making them the most reliable choices for practical deployment.
+In recent years, a real-time detection method based on DETR, known as RT-DETR [34], has also shown extraordinary performance.
+RT-DETR has also demonstrated extraordinary performance, achieving optimized inference speeds by designing an efficient encoder and simplifying its network architecture.
+In particular, YOLO11 primarily focuses on modifications in the head section, while its core feature extraction framework remains largely consistent with YOLOv8.
+Given the proven stability and mature migration performance of YOLOv8 in various scenarios, our work thus draws more on YOLOv8 for FENet design, ensuring reliability in our MOT pipeline, which prioritizes low latency over the transformer encoder.
 3 Method
 3.1 DecoderTracker
+3.1.1 Architecture of Proposed Methods
+The overall architecture of DecoderTracker and DecoderTracker+ are illustrated in the provided Figure 2.
+Each frame of the input video is fed into the Feature Extraction Network(FENet) to obtain multi-scale feature maps.
+For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is inputted to the decoder.
+For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder.
+These queries interact with the image features in the decoder to generate hidden states for bounding box predictions.
+Additionally, MOTR’s Track Association Network (TAN) is employed to update the track queries, providing trajectory queries for the next frame.
+3.1.2 Design of FENet
 Figure 3: The structure of FENet.
 It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences.
-Specific module details can be found in Figure 4.
-The extracted features will be fed into the Decoder and introduced to the cross-attention module to interact with the queries.
-Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder and the TAN.
-3.1.1 DecoderTracker Architecture
-DecoderTracker’s overall architecture is illustrated in the provided Figure 2.
-In this setup, a video sequence is input to image Feature Extraction Network(FENet) to extract features.
-For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
-For subsequent frames in the video sequence, the tracking queries from the previous frame and the learnable detect queries are concatenated and undergo filtering before being input to the decoder.
-These queries interact with the image features in the decoder to generate hidden states for bounding box predictions.
-Additionally, a feature aggregation process using MOTR’s TAN is performed to generate trajectory queries for the next frame.
-3.1.2 Design of FENet
-The structure of FENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction.
+The left part of the figure corresponds to the backbone, which consists of C2m and Conv3 modules (detailed in Figure 4) to extract initial multi-scale features from input images.
+The right part represents the neck, where features from the backbone are processed through upsampling and concatenation operations—specifically, high-resolution features from upper layers are fused with low-resolution features from lower layers via C2m modules—to generate the final three-scale feature maps.
+These feature maps are then fed into the decoder and introduced to the cross-attention module to interact with the queries.
+Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder.
+FENet is a lightweight backbone-plus-FPN network derived from YOLOv8.
+It replaces the ResNet-50 backbone and the multi-layer transformer encoder in MOTR, delivering three-scale feature maps to the decoder.
 The reason for not adopting the latest YOLOv9 or YOLOv10 is due to its slower training speed compared to YOLOv8.
-Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
-The design of this FENet is intended to enhance the method’s efficiency and increase its inference speed.
-Thus, our design primarily references the network structure of YOLOv8.
-The backbone structure of our method is almost identical to that of YOLOv8.
-Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
-Figure 4: Specific module details, modules not specified in the diagram are network structures or APIs that come with PyTorch.
+Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder along with the track and detect queries.
+Our FENet is intended to enhance the efficiency of the method and increase its inference speed.
+Thus, our design primarily resembles the YOLOv8 network structure.
+The only differences are in the neck structure, which we adapt to fit the input requirements of the decoder.
+Figure 4: Specific module details.
+Modules not specified in the diagram are network structures or APIs that come with PyTorch.
+The C2m module, composed of Split, BottleNeckF, and Concat operations, is a core component in both the backbone and neck of FENet (as shown in Figure 3), responsible for feature refinement and propagation.
+The SPPF module, located in the backbone, uses multiple MaxPool2d layers to enhance multi-scale feature aggregation.
 These modules are primarily constructed based on CNN, capable of efficiently extracting image features.
@@
 In the initial stage, the network is trained as a detection network, and the loss function is formulated defined as Eq.3:
-$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\mathrm{giou}}\mathcal{L}_{\mathrm{giou}}+\lambda_{enc}\mathcal{L}_{enc}$ (3)
-where $\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the frame predictions.
-The loss is calculated by ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$.
-$\mathcal{L}_{cls}$ denotes the focal loss[32], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[33].
-$\mathcal{L}_{enc}$ is the anchor loss of DAB-DETR [31] and $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients, which are set to 1,2,5,5 in our method, just like DAB-DETR.
-In this stage, the network is capable of initially learning the appearance features of objects to be tracked.
+$\mathcal{L}\left(\left.\widehat{Y_{t}}\right|_{\omega_{t}},Y_{t}\right)=\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{l_{1}}\mathcal{L}_{l_{1}}+\lambda_{\text{giou}}\mathcal{L}_{\text{giou}}+\lambda_{\text{enc}}\mathcal{L}_{\text{enc}}$ (3)
+where $\widehat{Y}=\{\widehat{Y}_{i}\}_{i=1}^{N}$ denotes frame predictions, $Y=\{Y_{i}\}_{i=1}^{N}$ is the ground truth, and $\omega=\{\omega_{i}\}_{i=1}^{N}$ represents matched indices.
+The component losses are defined as follows:
+$\mathcal{L}_{\text{cls}}$ $=\sum_{i=1}^{N}-\alpha\left(1-p_{i,c_{i}}\right)^{\gamma}\log\left(p_{i,c_{i}}\right)\quad$ (4)
+[…]
+$\mathcal{L}_{\text{giou}}$ $=\sum_{i=1}^{N}\left[1-\text{GIoU}\left(\hat{b}_{i},b_{i}\right)\right]\quad$ (6)
+$\mathcal{L}_{\text{enc}}$ $=\sum_{j=1}^{K}\left[\left\|\hat{b}^{\text{enc}}_{j}-b_{\sigma(j)}\right\|_{1}+\beta\left(1-\text{GIoU}\left(\hat{b}^{\text{enc}}_{j},b_{\sigma(j)}\right)\right)\right]\quad$ (7)
+Here, $p_{i,c_{i}}$ denotes the predicted probability for the ground truth class $c_{i}$ at the matched index $i$.
+The predicted and ground truth bounding boxes are represented by $\hat{b}_{i}$ and $b_{i}$, respectively, each defined by the center coordinates, width, and height.
+The $j$-th encoder-predicted anchor box is denoted as $\hat{b}^{\text{enc}}_{j}$.
+The matching process is governed by $\sigma$, a function that maps the top-$K$ encoder predictions to ground truth instances based on center distance, where $K=\min(\text{top-}K,N)$ is dynamically set to the number of ground truth objects.
+Hyperparameters include $\alpha=0.25$ and $\gamma=2$ for focal loss and $\beta=1$ for GIoU weight in anchor loss.
+The loss weights are configured as $\lambda_{\text{cls}}=1$, $\lambda_{l_{1}}=2$, $\lambda_{\text{giou}}=5$, and $\lambda_{\text{enc}}=5$ to balance classification, L1 regression, GIoU, and encoder losses.
+In addition, a more specific explanation of these losses can be found in [38, 39, 37].
+In this stage, the network is able to initially learn the appearance characteristics of the objects to be tracked.
 During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time.
-Additionally, various data augmentation techniques [34, 35, 36] can be employed to expedite the method’s extraction of appearance features.
-The second stage of training is conducted using MOTR’s Collective Average Loss.
+Furthermore, various data augmentation techniques [40, 41, 42] can be used to expedite the method’s extraction of appearance features.
+Whether these data augmentation methods are applied has a significant impact on the results of this stage, exerting an approximate one point effect on the final tracking accuracy.
+Since this is not the focus of this paper, detailed results are not enumerated here.
+The second stage of training is carried out using the Collective Average Loss of MOTR.
 During this phase, only MOTR’s TAN and our TBSP are utilized.
-The loss functions used in this stage and the third stage are as Eq.4.
+The loss functions used in this stage and the third stage are as in Eq.8.
 […]
@@
 At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects.
-Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage.
+Newly appeared objects in the second stage are defined as high-confidence predictions from detect queries (exceeding a confidence threshold) that pass the TBSP filtering process, i.e., their bounding boxes do not significantly overlap (IOU $\leq$ $\delta_{iou}$) with any existing track query (see Section 3.1.3 for details on TBSP).
+This ensures that they are distinct from already tracked objects.
+Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to that of the second stage.
 Thanks to our various strategies and improvements to the network, we achieved huge improvements in training time and inference speed compared to MOTR.
@@
 3.2.1 Fixed-Size Query Memory
-FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N\times d}$, where $N$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension.
+FSQM maintains fixed-dimensional queries $q_{tr}^{\mathcal{M}}\in\mathbb{R}^{N_{m}\times d}$, where $N_{m}$ denotes the preset maximum number of tracking targets and $d$ represents the feature dimension.
 Compared to dynamic query mechanisms, this approach introduces:
-$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N}}_{\text{confidencevector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (5)
+$\hat{\omega}_{tr}^{\mathcal{M}}=(\underbrace{q_{tr}^{\mathcal{M}}}_{\text{static queries}},\underbrace{C_{\text{tr}}\in\mathbb{R}^{N_{m}}}_{\text{confidence vector}},\underbrace{\mathcal{ID}}_{\text{global ID pool}},\underbrace{BBox^{\mathcal{M}}}_{\text{bounding box}})$ (9)
 to model memory for the continuous tracking process.
@@
 By the way, all inactive queries are assigned an ID of -1 and their bounding boxes are initialized with zero values.
+And the maximum number of tracking targets $N_{m}$ is set to be equal to the number of detect queries.
+Theoretically, as long as the actual number of tracked objects does not exceed this preset $N_{m}$, tracking performance remains unaffected.
+Based on empirical considerations, we thus set $N_{m}$ to match the number of object queries.
 During each frame processing, iterative updates are performed as follows:
@@
     $q_{tr}^{\mathcal{M}}[n]=\mathbf{0},\quad\mathcal{ID}[n]=-1.$ (12)
-    where $S(\cdot)$ represents the prediction score function. The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
-It is also important to emphasize that, aside from this, the processing of queries in FSQM is consistent with that in DecoderTracker.
-The specific process will not be repeated here, but can be referred to in the previous sections.
+    The decay ensures that the relevance of older queries decreases over time, while new, active queries retain higher confidence scores.
+It is also important to emphasize that, beyond the fixed-size query management (via injection/removal mechanisms) and the unique attention masking strategy of FSQM, the core handling of queries remains consistent with DecoderTracker: including the interaction between detect queries ($q_{d}$) and track queries ($q_{tr}$) in the decoder, cross-attention with FENet-derived feature maps, track query updates via TAN, and prediction of bounding boxes and object classes based on query hidden states.
 Algorithm 2 FSQM Online Update Procedure
@@
 6:        $q_{\text{tr}}[k]\leftarrow q_{j},BBox_{\text{tr}}[k]\leftarrow BBox_{j}$Momentum update}
-7:        $C_{\text{tr}}[k]\leftarrow c_{i}$Reset confidence to 1.0}
+7:        $C_{\text{tr}}[k]\leftarrow 1$Reset confidence to 1.0}
 8:     end for
@@
 16:  return $\mathcal{M}^{\prime}=(Q_{\text{tr}},C_{\text{tr}},\mathcal{ID})$
-3.2.2 Attention Mechanism Optimization
+3.2.2 Attention Masking Strategy
 Since some queries in FSQM are empty, we need to explore the network structure to ensure that these empty queries do not affect the final results.
@@
 4.1 Datasets and Metrics
-To validate DecoderTracker and DecoderTracker+’s performance, we evaluate the algorithm on three demanding datasets: DanceTrack[37], MOT17[38], and KITTI[39].
+Datasets: To validate the performance of DecoderTracker and DecoderTracker +, we evaluated the method on three demanding datasets: DanceTrack[43], MOT17[44], and KITTI[45].
 DanceTrack is a vast dataset designed for human tracking, showcasing scenarios with occlusion, frequent crossovers, uniform appearances, and varied body gestures.
 Comprising 100 videos featuring diverse dance styles, it underscores the significance of motion analysis in multi-object tracking.
-MOT17 focuses on multiple object tracking in public spaces, primarily pedestrians.
+MOT17 focuses on the tracking of multiple objects in public spaces, primarily by pedestrians.
 The dataset includes seven scenes, divided into training and testing clips, providing object detection annotations suitable for both online and offline tracking approaches.
-KITTI serves as a research dataset for autonomous driving, encompassing stereo, optical flow, odometry, 3D object detection, and tracking tasks.
-Captured across diverse environments using advanced sensors, it provides raw data and task-specific benchmarks.
+KITTI covers stereo, flow, odometry, detection and tracking; we use its multi-object tracking split for evaluation.
 Evaluation Metrics: We assessed our method using widely recognized MOT evaluation metrics.
-The primary metrics employed for evaluation included HOTA, AssA, DetA ,IDF1 and MOTA[40, 41].
+The primary metrics employed for evaluation include HOTA, AssA, DetA ,IDF1 and MOTA[46, 47].
 These metrics collectively provided a comprehensive and robust evaluation of our method’s tracking performance.
@@
 Additionally, to better contrast with the MOTR series in terms of speed, inference speed tests were carried out on a single NVIDIA V100 GPU.
-One more thing, the image size for the input network is $640\times 640$.
+Furthermore, the image size for the input network is $640\times 640$.
+The number of object queries is configured as 60 for DanceTrack and KITTI and 300 for MOT17.
+These values are in line with the previous works [35, 22].
+Consistently, the preset maximum number of tracking targets in the FSQM system is also aligned with these query numbers.
 (a) Comparison of Training Time on MOT17
@@
 The training time of DecoderTracker+ is basically the same as that of DecoderTracker, so it will not be listed separately here.
-Figure 7: IDF1-MOTA-HOTA and IDF1-MOTA-FPS comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on Dancetrack test sets.
+Figure 7: IDF1-MOTA-HOTA and IDF1-FPS-HOTA comparisons of state-of-the-art trackers with our proposed DecoderTracker and DecoderTracker+ on DanceTrack test sets.
 For the left image, the horizontal axis is MOTA, the vertical axis is IDF1, and the radius of the circle is HOTA.
@@
 Our DecoderTracker+* achieves the best IDF1 and HOTA and comparable speed performance.
-Table 2: Performance comparison between DecoderTracker and existing methods on the Dancetrack[37] test set.
-Numbers are highlighted in bold when they correspond to either DecoderTracker or the superior metric from MOTR.
-* means that extra data was used during training.
-Methods IDF1 HOTA AssA DetA MOTA
-Non-End-to-end:
-TransTrack[15] 45.2 45.5 27.5 75.9 87.7
-GTR[21] 50.3 48.0 31.9 72.5 84.7
-DST-track[42] 51.0 51.9 34.6 82.3 84.9
-ByteTrack[14] 52.5 47.3 31.3 71.6 89.5
-OC-SORT[22] 54.2 54.6 38.0 80.4 89.4
-Deep OC-SORT[5] 61.5 61.3 45.8 82.2 92.3
-End-to-end:
-MOTR[2] 51.5 54.2 40.2 73.5 79.7
-DecoderTracker 54.2 54.2 35.7 78.7 87.8
-DecoderTracker+ 54.4 54.5 35.6 79.0 88.1
-DecoderTracker* 62.9 60.8 47.1 81.2 91.1
-DecoderTracker+* 62.2 61.2 46.0 81.8 91.9
-Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [38] set. * means that extra data was used during training.
-method IDF1 HOTA DetA AssA MOTA
-Non-End-to-end:
-YOLOTracker[43] 65.1 - - - 71.9
-FairMOT[44] 72.3 59.3 60.9 58.0 73.7
-MSPNet[45] 74.6 - - - 71.9
-ByteTrack[14] 77.2 62.8 63.8 62.2 78.9
-OC-SORT[22] 77.5 63.2 63.2 63.2 78.0
-StrongSORT[21] 78.3 63.5 63.6 63.7 78.5
-UTM[46] 78.7 64.0 65.9 62.5 81.8
-Deep OC-SORT[5] 80.6 64.9 64.1 65.9 79.4
-End-to-end:
-MOTR[2] 59.4 57.8 60.3 55.7 73.4
-DecoderTracker 70.5 55.7 55.7 56.0 66.5
-DecoderTracker+ 70.4 55.8 55.2 55.8 66.7
-DecoderTracker* 77.2 61.8 62.1 60.1 76.8
-DecoderTracker+* 77.5 62.2 62.5 60.7 76.7
-During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti. Subsequent stages utilized a uniform batch size of 1.
-For MOT17, like MOTR, we integrated extra data from the CrowdedHuman [47], employing random displacement to generate video segments with pseudo-trajectories.
+During the initial training, batch sizes were set to 32 on the 4090 and 16 on the 2080ti.
+Subsequent stages utilized a uniform batch size of 1.
+The random seed was fixed at 42 throughout all experiments to ensure reproducibility.
+For MOT17, following MOTR’s data augmentation strategy [2], we incorporated additional data from CrowdHuman [48].
+To convert static images into video-like sequences, we applied random affine transformations (primarily random translation and rotation) to individual images across simulated frames.
 Training comprised 120 epochs in the first stage, followed by 30 epochs in the second.
@@
 The phased training regimen—12.1 hours for feature embedding, 30.8 hours for temporal modeling, and 43.4 hours for TBSP-enhanced prediction—systematically constructs tracking capabilities while maintaining 83% shorter wall-clock time than competitors.
-Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS,
-For DanceTrack ,the experimental results shown in Table 2 reveal that DecoderTracker and its enhanced variant DecoderTracker+ achieve comparable precision levels across core evaluation metrics.
-Both models demonstrate superior detection accuracy over existing end-to-end methods, with DetA scores of 78.7 and 79.0 respectively, significantly outperforming MOTR’s 73.5.
-Their MOTA metrics (87.8 and 88.1) approach SOTA non-end-to-end methods like ByteTrack (89.5), while maintaining balanced HOTA performance at 54.2 and 54.5.
-This parity suggests that architectural refinements between the two variants primarily yield marginal improvements rather than fundamental performance shifts.
-Notably, both configurations exhibit identical IDF1 scores (54.2 and 54.4), confirming consistent identity preservation capabilities.
-The minimal performance delta between variants implies optimized parameter utilization rather than architectural overhauls.
-When extended with supplementary training data (DecoderTracker* variants), both models achieve substantial gains, attaining 91.1-91.9 MOTA and 60.8-61.2 HOTA.
-While their AssA scores (35.6-35.7) remain below MOTR’s 40.2, the detection-centric advantages manifest through DetA superiority over all end-to-end and most non-end-to-end competitors.
-These findings position the DecoderTracker framework as a robust baseline for end-to-end tracking, achieving detection reliability comparable to sophisticated multi-stage systems while preserving model efficiency.
-In the MOT17 benchmark (Table 3), DecoderTracker shows strong performance, especially in complex tracking scenarios.
-While slightly behind MOTR in HOTA and AssA, it compensates with comparable DetA and excels in IDF1 by nearly 10 points.
-This advantage is due to its simplified network structure, which achieves competitive accuracy with less training data.
-Moreover, DecoderTracker maintains robust overall performance, as shown by its high MOTA score, making it suitable for real-world applications.
-When additional training data is introduced, DecoderTracker’s performance further improves.
-For instance, DecoderTracker* achieves a significant boost in IDF1 to 77.2, HOTA to 61.8, and MOTA to 76.8.
-Similarly, DecoderTracker+* also shows substantial gains.
-This clearly demonstrates the positive impact of more training data on DecoderTracker’s tracking capabilities, highlighting its potential for even better performance with increased data support.
-Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI [39] set.
+Inference performance reveals more pronounced advantages: DecoderTracker+ attains 28.8 FPS, 4.1× faster than MOTRv2’s 6.9 FPS.
+Table 2: Performance comparison between DecoderTracker and existing methods on the DanceTrack [43] test set.
+Bold numbers indicate either DecoderTracker results or superior MOTR metrics, while our method uses color highlighting for distinction.
+Asterisk (*) denotes methods using extra training data.
+Methods Metrics
+IDF1 HOTA AssA DetA MOTA
+Non-End-to-end
+TransTrack [17] 45.2 45.5 27.5 75.9 87.7
+ByteTrack [16] 52.5 47.3 31.3 71.6 89.5
+OC-SORT [27] 54.2 54.6 38.0 80.4 89.4
+MotionTrack [5] 58.6 58.2 41.7 81.4 91.3
+Deep OC-SORT [6] 61.5 61.3 45.8 82.2 92.3
+End-to-end
+DNMOT [25] 49.7 53.5 87.7
+MOTR [2] 51.5 54.2 40.2 73.5 79.7
+DecoderTracker 54.2 54.2 35.7 78.7 87.8
+DecoderTracker+ 54.4 54.5 35.6 79.0 88.1
+DecoderTracker* 62.9 60.8 47.1 81.2 91.1
+DecoderTracker+* 62.2 61.2 46.0 81.8 91.9
+As evidenced in Table 2, DecoderTracker establishes new performance benchmarks for end-to-end multi-object tracking on the challenging DanceTrack benchmark.
+Our method demonstrates superior detection capabilities, with DecoderTracker and DecoderTracker+ achieving substantially higher DetA scores than all existing end-to-end approaches.
+This detection advantage directly translates to competitive MOTA performance that approaches state-of-the-art non-end-to-end methods, underscoring the efficacy of our detection-centric architecture.
+Notably, both variants maintain balanced overall tracking performance comparable to the strongest end-to-end baseline while exhibiting consistent identity preservation metrics (IDF1).
+The minimal performance variance between DecoderTracker and its enhanced version suggests optimized architectural efficiency rather than fundamental design differences.
+When augmented with additional training data, our framework achieves transformative performance gains, matching or exceeding leading non-end-to-end methods across multiple metrics.
+These data-enhanced variants particularly showcase exceptional MOTA and IDF1 results, demonstrating our architecture’s scalability and untapped potential.
+While association accuracy (AssA) remains an area for future improvement relative to specialized baselines, the consistent DetA superiority highlights our method’s core strength in target localization.
+This positions DecoderTracker as a robust foundational framework that bridges the performance gap between end-to-end paradigms and complex multistage trackers while maintaining model efficiency.
+Table 3: Performance comparison between DecoderTracker and existing methods on the MOT17 [44] test set.
+Bold numbers indicate either DecoderTracker results or superior MOTR metrics, while our method uses color highlighting for distinction.
+Asterisk (*) denotes methods using extra training data.
+Methods Metrics
+IDF1 HOTA DetA AssA MOTA
+Non-End-to-end
+MSPNet [30] 74.6 71.9
+ByteTrack [16] 77.2 62.8 63.8 62.2 78.9
+OC-SORT [27] 77.5 63.2 63.2 63.2 78.0
+TWIX [29] 76.2 63.1 62.1 63.1 78.1
+Deep OC-SORT [6] 80.6 64.9 64.1 65.9 79.4
+End-to-end
+MOTR [2] 68.9 57.8 60.3 55.7 73.4
+DNMOT [25] 68.1 58.0 75.6
+MeMOT [50] 69.0 56.9 55.2 72.5
+DecoderTracker 70.5 55.7 55.7 56.0 66.5
+DecoderTracker 70.5 55.7 55.7 56.0 66.5
+DecoderTracker+ 70.4 55.8 55.2 55.8 66.7
+DQFormer [26] 71.6 59.2 60.6 58.1 75.2
+DecoderTracker* 73.3 61.7 62.5 60.4 76.7
+DecoderTracker+* 73.4 62.0 62.1 60.2 77.2
+In the MOT17 benchmark (Table 3), DecoderTracker exhibits strong performance, with a distinct advantage in IDF1 despite marginal deficits relative to MOTR in HOTA and AssA.
+Its simplified architecture achieves competitive accuracy with reduced training data while maintaining robust overall performance (high MOTA) suitable for real-world applications.
+Augmented training data further enhances its capabilities: DecoderTracker* and DecoderTracker+* demonstrate significant improvements in IDF1, HOTA, and MOTA, underscoring the positive impact of increased data volume.
+The slight gaps in HOTA and AssA compared to MOTR can be derived from two factors: first, the streamlined architecture (encoder removed) affects the modeling of long-term temporal correlations, which is critical for HOTA and AssA, which are sensitive to inter-object association accuracy; second, the weak supervision based on TBSP in the three-stage training strategy constrains the precision of the association in complex scenarios, in contrast to the more aggressive TALA strategy of MOTR.
+To address this, we propose three potential future improvements: integrating the release-fetch supervision from MOTRv3 and the coopetition label assignment strategy from CO-MOT to balance the label assignment weights between detection and tracking queries; enhancing the temporal modeling capability of the TAN module by introducing a dynamic adjustment mechanism for cross-frame attention; and optimizing the matching strategy between static queries and dynamic objects in FSQM to reduce interference from invalid queries on association accuracy.
+It should be noted that since the focus of this paper is to verify the efficiency advantages of the decoder-only architecture and the effectiveness of the basic training strategy, the improvement directions above will be explored in subsequent studies.
+Due to the constraints in length and scope of the current research, in-depth exploration of these directions is not included in this paper.
+Table 4: Performance comparison between DecoderTracker and existing methods on the KITTI [45] test set. Bold numbers indicate the best performance in each category.
 Pedestrian Car
-Tracker HOTA MOTA AssA HOTA MOTA AssA
-Non-End-to-end:
+Methods HOTA MOTA AssA HOTA MOTA AssA
+Non-End-to-end
 CenterTr [51] 40.3 53.8 36.9 73.0 88.8 71.1
@@
 TripletTrack [54] 50.0 46.5 57.8 73.6 84.3 74.7
-End-to-end:
+End-to-end
 DecoderTracker 51.5 56.8 58.4 72.2 83.2 73.8
@@
 Table 4 provides an overview of DecoderTracker’s performance on the KITTI benchmark.
-DecoderTracker achieves competitive results, with an HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively.
+DecoderTracker achieves competitive results, with HOTA of 70.2 and 82.9 for Car tracking and Pedestrian tracking, respectively.
 These results position DecoderTracker as a strong contender in MOT on the KITTI.
@@
 4.4.1 More Discusstion About TBSP
-The initial reason for developing the DecoderTrack method was not due to slow inference speed, but rather the extensive training resources required.
+The initial reason for developing the DecoderTracker method was not due to slow inference speed, but rather the extensive training resources required.
 Using eight V100 GPUs with pretrained weights, training on an extended MOT17 dataset, which contained approximately 10,000 images, required about 84 hours for 200 epochs.
@@
 Typically, these methods are pretrained on unlabeled data to learn image features, then fine-tuned on strongly supervised tasks such as classification or detection.
-We believe that the TBSP, a weak supervision process, enables the method to preliminarily learn object motion information that would otherwise require many more training iterations.
-It’s akin to learning to ride a bicycle where training wheels help beginners to ride without obstacles while also giving them a feel for biking, which aids in eventually mastering the skill.
-A similar situation may exist here.
 Although TBSP did not enhance method performance, it reduced the number of iterations needed for convergence, which might be considered its greatest contribution.
@@
 In YOLOv8, there is a BottleNeck module that differs slightly between the neck and backbone; we eliminated this difference and replaced the 3x3 convolution in the middle with a 1x1 convolution to further reduce the computational load.
-These modifications are illustrated in Figure 9, and their impact on performance is shown in Table 8.
+These modifications, grounded in YOLOv8’s proven stability and migration performance (with YOLO11’s key changes focusing more on the head section, less relevant to our feature extraction adjustments), are illustrated in Figure 9, and their impact on performance is shown in Table 8.
 Figure 9: The difference between our proposed method and YOLOv8 in terms of modules.
@@
 YOLOv8 51.4 52.4 84.2 36.1 18.9
-1.88 DecoderTracker+:
+DecoderTracker+:
 FENet 53.6 53.4 84.9 35.7 28.2
@@
 Figure 10: DecoderTracker and MOTR failed in the same scenario. The probable cause may lie in the insufficient scale of the MOT17 benchmark.
-This paper proposes an efficient decoder-only end-to-end MOT method, which achieves faster inference speed and shorter training time compared to the baseline MOTR.
-Additionally, we introduce new training strategies aimed at accelerating the convergence of the method.
-Due to hardware limitations, we were unable to verify our training strategies on MOTR, but we believe these strategies are generalizable.
-Finally, we conducted a detailed and in-depth analysis of the slow inference process in MOTR, finding that dynamic data significantly impacts inference speed.
-Subsequently, we applied FSQM to address this issue, greatly enhancing the inference speed of the tracking-by-query end-to-end paradigm.
-Overall, this study provides a more efficient benchmark method for the multi-object tracking field, lays the foundation for the emergence of real-time end-to-end MOT methods, and offers insights for future development.
-In future work, we should further explore more efficient network architectures to build real-time MOT methods.
-Of course, we also plan to integrate the technologies from MOTRv2, MOTRv3, and CO-MOT into DecoderTracker to improve tracking performance.
-Undoubtedly, the greatest limitation of DecoderTracker, similar to that of MOTR, is its performance on the MOT17 benchmark, which does not match that of tracking-by-detection methods.
-Both methods fail under similar circumstances, as illustrated in Figure.10.
-This issue may, like with MOTR, stem from the insufficient data volume on the MOT17 benchmark, which does not adequately train the method.
-Clearly, despite this limitation, DecoderTracker still represents a significant improvement over MOTR.
-Although we attempted to use a mixed dataset to address this issue, our method still underperforms compared to SOTA tracking-by-detection methods.
-How to resolve this issue will also be a key focus of future research.
+This paper presents DecoderTracker and its enhanced variant DecoderTracker+, two novel end-to-end MOT methods based on a decoder-only architecture, aiming to address the issues of high computational resource consumption, suboptimal inference speed, and prolonged training time in traditional transformer-based end-to-end MOT methods like MOTR.
+DecoderTracker optimizes the network architecture by removing the redundant encoder and adopting a lightweight Feature Extraction Network (FENet) derived from YOLOv8, which enhances efficiency and inference speed.
+It also introduces a three-stage training strategy, incorporating TBSP as a weakly supervised training step to accelerate convergence and reduce training time.
+To tackle optimization challenges from dynamic data, DecoderTracker+ is proposed by integrating FSQM and refining attention layers with an attention masking strategy.
+FSQM maintains a constant number of tracking queries, enabling better compatibility with modern deep learning frameworks and facilitating deployment, while the masking strategy ensures that inactive queries do not interfere with computations.
+Experimental results on benchmarks including DanceTrack, MOT17, and KITTI demonstrate that both methods outperform MOTR without additional tricks.
+DecoderTracker+ achieves an inference speed of 28.8 FPS, which is 2-3 times faster than MOTR, while maintaining competitive performance in metrics such as HOTA, IDF1, and MOTA.
+This work provides an efficient baseline for end-to-end MOT, laying the foundation for real-time end-to-end MOT methods and offering insights for engineering deployment, with potential future improvements including enhancing temporal modeling and optimizing label assignment strategies .
+DecoderTracker’s primary limitation remains its accuracy gap on MOT17 compared to TBD(e.g., Deep OC-SORT), similar to MOTR’s challenge.
+We attribute this in part to the limited volume of MOT17 training data, which affects end-to-end training effectiveness (Figure 10).
+Although our method significantly improves speed and outperforms MOTR, it still trades accuracy for real-time efficiency, a core design choice.
+Additional challenges include robustness in high-occlusion scenarios (Fig.
+10) and domain generalization.
+Future work will focus on improving occlusion handling, exploring domain adaptation, and refining label assignment strategies to close the accuracy gap.